### v3

Abstract
In this paper, we propose a novel kernel stochastic gradient descent (SGD) algorithm for large-scale supervised learning with general losses. Compared to traditional kernel SGD, our algorithm improves efficiency and scalability through an innovative regularization strategy. By leveraging the infinite series expansion of spherical radial basis functions, this strategy projects the stochastic gradient onto a finite-dimensional hypothesis space, which is adaptively scaled according to the bias-variance trade-off, thereby enhancing generalization performance. Based on a new estimation of the spectral structure of the kernel-induced covariance operator, we develop an analytical framework that unifies optimization and generalization analyses. We prove that both the last iterate and the suffix average converge at minimax-optimal rates, and we further establish optimal strong convergence in the reproducing kernel Hilbert space. Our framework accommodates a broad class of classical loss functions, including least-squares, Huber, and logistic losses. Moreover, the proposed algorithm significantly reduces computational complexity and achieves optimal storage complexity by incorporating coordinate-wise updates from linear SGD, thereby avoiding the costly pairwise operations typical of kernel SGD and enabling efficient processing of streaming data. Finally, extensive numerical experiments demonstrate the efficiency of our approach.
Keywords and phrases: Kernel stochastic gradient descent; Online learning; General losses; Spherical radial basis functions; Optimal convergence.
Mathematics Subject Classification (2020): 68T05, 68Q32, 62L20
1 Introduction
[…]
$\min_{f\in\mathcal{W}}\mathcal{E}(f):=\min_{f\in\mathcal{W}}\mathbb{E}_{\rho}\left[\ell(f(X),Y)\right],$ (1.1)
where $\mathcal{W}$ is a subset of an infinite-dimensional reproducing kernel Hilbert space (RKHS) induced by a kernel $K(x,x^{\prime})$ constructed from spherical radial basis functions (see Subsection 2.2 for details).
In kernel-based algorithms, appropriate regularization strategies play a crucial role in enhancing generalization performance. Traditional kernel-based stochastic gradient descent (SGD) typically introduces regularization by approximating the regularization path or adjusting the step size. However, these approaches are not imposed directly on the hypothesis space and therefore have only a limited influence on its complexity. As a result, the hypothesis space in traditional kernel SGD does not adapt to the difficulty or ill-conditioning of the problem (1.1), which may cause excessively rapid variance accumulation and lead to suboptimal convergence rates. In contrast, the stochastic approximation framework proposed in this paper updates the estimator by projecting $K(X_{n},\cdot)$ onto a finite-dimensional hypothesis space tailored to the difficulty of the problem. We show that this regularization strategy not only improves generalization but also substantially reduces computational complexity, while ensuring optimal storage complexity. Specifically, the algorithm requires only $\mathcal{O}(n^{1+\frac{d}{d-1}\epsilon})$ time and $\mathcal{O}(n^{\epsilon})$ memory, where $n$ denotes the sample size. The parameter $\epsilon\in\left(0,\frac{1}{2}\right)$ can be chosen arbitrarily small, provided that the minimizer or the underlying hypothesis space possesses sufficient smoothness.
1.1 Related Works and Discussion
Nonparametric regression based on reproducing kernels is both theoretically well understood and widely applied across diverse areas of science and engineering [51, 59, 43, 61, 55, 39, 62]. Recent work has investigated the comparability between specific classes of deep neural networks and kernel methods [29, 67], sparking growing interest in scalable kernel techniques for large datasets. Within the framework of nonparametric least-squares regression under batch learning—where the entire dataset is available upfront—substantial progress has been made toward improving the computational efficiency of large-scale kernel methods [45, 66, 3, 62, 48, 1]. Algorithms such as EigenPro 3.0 [1] and FALKON-BLESS [48] leverage gradient-based optimization, preconditioning strategies, and low-rank kernel approximations to effectively reduce both storage requirements and computational costs. The quadratic structure of the least-squares loss, in particular, greatly simplifies theoretical analysis and facilitates practical implementation [66, 49]. Despite these advantages, the lack of Lipschitz continuity in the least-squares loss makes the estimator highly sensitive to outliers. From a robustness perspective, non-quadratic losses, such as the Huber loss and the logistic loss, are often preferred. Consequently, earlier works [65, 7, 14] studied the statistical properties of such losses, including consistency and robustness, while more recent studies [36, 2, 57, 60] analyze the convergence of empirical risk minimization (ERM) with non-quadratic losses. However, efficient optimization with these losses on large-scale datasets remains a significant barrier. Unlike least-squares loss, where regularized ERM admits closed-form solutions, non-quadratic losses typically lack explicit expressions and instead require iterative numerical solvers, thereby incurring additional computational costs. Existing large-scale kernel methods are primarily designed for least-squares loss, and extending them to handle non-quadratic losses without sacrificing efficiency is a nontrivial task. Designing a kernel method that is both computationally scalable and statistically optimal for general losses, thus remains an open and pressing problem.
In online learning, where samples arrive sequentially, the estimator must be updated upon receiving each sample. This naturally motivates the use of SGD, known for its efficiency in optimization [46, 44, 33, 12, 30]. Consequently, SGD has been widely applied to nonparametric least-squares regression, giving rise to kernel SGD [32]. A series of studies have analyzed the convergence of kernel SGD, beginning with [53, 63], and subsequently refined in [19] toward achieving optimal rates in [24, 64]. More specifically, the difficulty of the nonparametric least-squares regression problem is characterized by the spectral structure of the Hessian and by the regularity conditions that describe the smoothness of the optimal solution. Since the least-squares loss and related risk functionals (e.g., population risk, excess risk), which measure the generalization performance of the algorithm, are quadratic, the gradient of these risks reduces to an analytically tractable linear operator. As a result, convergence analyses in this setting typically rely on precise characterizations of the Hessian operator and the associated trace inequalities. In contrast, analyzing general loss functions is considerably more challenging: the Hessian of the population risk (1.1) is generally a nonlinear operator depending on $f\in\mathcal{W}$, unlike in the least-squares case, where its Hessian simplifies to a fixed and well-understood covariance operator independent of $f$. In such cases, analyzing the properties of the Hessian operator, particularly precisely characterizing its spectral structure, is highly nontrivial. From an optimization perspective, (1.1) can be reformulated as a stochastic optimization problem with ill-conditioned objectives, since the Hessian eigenvalues typically decay to zero. For ill-conditioned instances of (1.1), classical optimization techniques—typically applicable in finite-dimensional hypothesis spaces and without requiring regularity of the optimal solution—yield at usual optimal slow rate $\mathcal{O}\left(\tfrac{1}{\sqrt{n}}\right)$ [52]. However, if the objective function is well-conditioned (i.e., the eigenvalues of the Hessian are bounded away from zero), SGD in finite-dimensional spaces generally attains the optimal rate $\mathcal{O}\left(\tfrac{1}{n}\right)$ [4, 52]. In the case of nonparametric least-squares regression in infinite-dimensional hypothesis spaces, strong regularity conditions on the optimal solution can improve the well-posedness of (1.1), thereby enabling convergence rates faster than $\mathcal{O}\left(\tfrac{1}{\sqrt{n}}\right)$. This motivates us to integrate optimization techniques with generalization analysis under regularity assumptions, with the goal of establishing fast convergence rates for kernel SGD with general losses, in analogy to the least-squares setting.
In the online setting, although the generalization performance of kernel SGD has been extensively investigated, it inevitably incurs a quadratic computational cost in the sample size [56, 19], since each update requires operations over all pairs of samples. In our recent work [5], we proposed a kernel SGD algorithm for the least-squares loss that incorporates coordinate-wise updates, inspired by linear SGD [^1] Linear SGD is equivalent to kernel SGD with a linear kernel [16].. Compared with standard kernel SGD, this algorithm not only reduces the computational burden but also overcomes the saturation phenomenon in convergence rates—a limitation widely observed in the analysis of kernel SGD [19, 24]—thereby achieving statistical optimality. Numerical experiments further show that, relative to popular large-scale kernel methods in the batch setting [19, 49, 50, 1], the proposed algorithm delivers superior empirical performance, exhibiting faster convergence of the generalization error with comparable running time. Building on this foundation, the present paper introduces a novel kernel SGD framework for general losses that preserves both computational efficiency and statistical optimality.
1.2 Algorithm Overview and Main Contributions
Based on spherical radial basis functions (SBFs), we propose a novel SGD algorithm for general losses. The underlying hypothesis space $\mathcal{H}$ is an infinite-dimensional RKHS induced by SBFs, which naturally incorporates the geometry of the spherical manifold. Exploiting the infinite series expansion of SBFs, we construct an increasing sequence of finite-dimensional nested subspaces $\{\mathcal{H}_{L_{n}}\}_{n\geq 0}\subset\mathcal{H}$, where $\mathcal{H}_{L_{n}}$ serves as the hypothesis space at the $n$-th iteration of SGD. Specifically, upon receiving the $n$-th sample, the estimator is updated along the negative direction of the projection of the stochastic gradient of (1.1) onto $\mathcal{H}_{L_{n}}$. This amounts to truncating the original gradient within $\mathcal{H}_{L_{n}}$, and we therefore refer to this approach as truncated kernel stochastic gradient descent, or T-kernel SGD for short. As samples arrive sequentially, the algorithm adaptively tunes its regularization strength by controlling the complexity of the hypothesis space $\mathcal{H}_{L_{n}}$. In Section 2, we show that the projected stochastic gradient onto $\mathcal{H}_{L_{n}}$ admits an explicit closed-form expression. For the output, we adopt suffix averaging [52], which combines the advantages of Polyak averaging and the last iterate, thereby enhancing robustness and accelerating convergence. Moreover, by constructing a $C^{1}$-diffeomorphism $F$ between a general closed domain $\Omega$ and $\mathbb{S}^{d-1}$, we extend T-kernel SGD originally designed for spherical inputs to arbitrary input domains $\Omega$. Our convergence analysis is thus developed in general spaces, ensuring broad applicability beyond spherical data. From a technical perspective, we characterize the spectral structure of the covariance operator and the regularity conditions via two sequences of norm-based asymptotic inequalities. Together with tools from stochastic optimization in Hilbert spaces, this allows us to establish convergence guarantees without relying heavily on Hessian operators. Building on this framework, we prove that T-kernel SGD achieves minimax optimal rates for general losses, up to a logarithmic factor. Furthermore, we establish an optimal strong convergence result in RKHS, which, to the best of our knowledge, is novel in the context of general losses. Such convergence typically implies uniform convergence of higher-order derivatives [54], yet has rarely been studied in the literature. Finally, when the minimizer of (1.1) exhibits sufficient smoothness, T-kernel SGD achieves computational and optimal memory complexities: $\mathcal{O}(n^{1+\frac{d}{d-1}\epsilon})$ in time and $\mathcal{O}(n^{\epsilon})$ in memory, where $0<\epsilon<\tfrac{1}{2}$ can be arbitrarily small.
The remainder of the paper is structured as follows. In Section 2, we introduce the basic assumptions on loss functions and briefly review the theoretical background of SBFs. We then propose the T-kernel SGD with general losses. We introduce the mathematical framework underlying T-kernel SGD and present its convergence behaviors in Section 3. In Section 4, we validate the theoretical guarantees and analyze the computational complexity through numerical experiments [^2] The code for the experimental section of the paper is located at the link: https://github.com/Researcher-Bai-1/T-kernel-SGD-with-general-losses.git. All proofs of the theorems are deferred to the Appendix.
2 Preliminaries and Algorithm
In this section, we outline the basic assumptions on the loss functions and give examples that satisfy them. We then review the theoretical foundations of spherical radial basis functions and their role in defining the hypothesis space. Finally, we introduce truncated kernel stochastic gradient descent and discuss its extension to broader input domains.
2.1 Loss Functions
The primary objective of this paper is to infer the function $f^{*}$ that minimizes the population risk over a subset $\mathcal{W}$ of the underlying hypothesis space, i.e.,
$f^{*}:=\arg\min_{f\in\mathcal{W}}\mathcal{E}(f)=\arg\min_{f\in\mathcal{W}}\mathbb{E}_{\rho}\left[\ell\left(f\left(X\right),Y\right)\right]$
where $\ell(u,v):\mathbb{R}\times\mathcal{Y}\to\mathbb{R}_{+}$ denotes a loss function. In this subsection, we state the assumptions that characterize the loss function $\ell(u,v)$. Intuitively, when the loss function exhibits locally quadratic behaviour, one may expect the algorithm to achieve the same convergence rate as that obtained with the least-squares loss. At the same time, our assumptions are broad enough to encompass many standard losses in supervised learning, including least-squares, logistic, Poisson, and Cauchy losses. We next introduce several commonly used notions, such as local strong convexity and local smoothness. Throughout the paper, we restrict our attention to the domain $B,B]\times\mathcal{Y}$, where $B>0$ is a fixed constant.
Assumption 1.
On the domain $B,B]\times\mathcal{Y}$, the loss function $\ell(u,v)$ is partially differentiable with respect to $u$, and its derivative is uniformly bounded; that is, there exists a constant $M>0$ such that $|\partial_{u}\ell(u,v)|\leq M$ for all $(u,v)\in[-B,B]\times\mathcal{Y}$.
Assumption 2.
(Local $L$-smoothness) The loss function $\ell(u,v)$ is locally $L$-smooth on [B,B][-B,B]; that is, there exists a constant $L>0$ such that for all $u_{1},u_{2}\in[-B,B]$, it holds
$\left|\partial_{u}\ell(u_{1},v)-\partial_{u}\ell(u_{2},v)\right|\leq L|u_{1}-u_{2}|,\ \ \forall v\in\mathcal{Y}.$ (2.1)
Assumption 3.
(Local $\mu$-strong convexity) The loss function $\ell(u,v)$ is locally $\mu$-strongly convex with respect to its first argument $u$ over the interval [B,B][-B,B]; that is, there exists a constant $\mu>0$ such that for all $u_{1},u_{2}\in[-B,B]$, one has
$\ell(u_{1},v)-\ell(u_{2},v)-\partial_{u}\ell(u_{2},v)(u_{1}-u_{2})\geq\frac{\mu}{2}(u_{1}-u_{2})^{2},\ \ \forall v\in\mathcal{Y}.$ (2.2)
Assumption 1 and Assumption 2 together guarantee the existence of the Fréchet derivative (see, e.g., [15]) of the population risk, thereby ensuring that the stochastic gradient descent algorithm is well-defined. Local smoothness, as formalized in Assumption 2, is a standard and widely adopted assumption in the optimization [41]. In finite-dimensional hypothesis spaces, the locally strong convexity of the loss is sufficient to guarantee the optimal rate $\mathcal{O}(\frac{1}{n})$ [4, 30], while assuming only convexity typically leads to the slow rate $\mathcal{O}(\frac{1}{\sqrt{n}})$ [52]. Assumption 2 and Assumption 3 are essential for establishing the fast rates we aim to prove. Moreover, these assumptions can be readily verified under the following sufficient condition: if the second-order partial derivative $\partial_{uu}^{2}\ell(u,v)$ is positive and bounded above by $L>0$ and below by $\mu>0$ on $B,B]\times\mathcal{Y}$, and if $\partial_{u}\ell(u,v)$ is also bounded, then Assumption 1, Assumption 2, and Assumption 3 hold.
In nonparametric regression, the output space $\mathcal{Y}$ is typically assumed to be a subset of $\mathbb{R}$ [55, 56, 19]. By contrast, in our framework $\mathcal{Y}$ may be any nonempty set, allowing the response variable $Y$ to take values in a discrete set for classification or to represent sequences, functional data, and other types of outputs. Furthermore, our framework accommodates globally non-convex loss functions, including the Cauchy loss [9] and the Welsch loss [27]. Below, we list several commonly used losses in supervised learning that satisfy our assumptions. Unless otherwise specified, $B$ denotes an arbitrary fixed positive real number in the examples that follow.
  • Least-square loss: $\ell(u,v)=(u-v)^{2}$, where $(u,v)\in[-B,B]\times\mathcal{Y}$ and $\mathcal{Y}$ is a compact subset of $\mathbb{R}$.
  • Logistic loss: $\ell(u,v)=\log(1+e^{-vu})$, where $(u,v)\in[-B,B]\times\mathcal{Y}$ and $\mathcal{Y}=\{-1,1\}$.
  • Loss in Poisson regression [21]: $\ell(u,v)=e^{u}-uv$, where $(u,v)\in[-B,B]\times\mathcal{Y}$ and $\mathcal{Y}$ is a finite set in $\mathbb{N}$}.
  • Huber loss [25]: $\ell(u,v)=W(v-u)$, for $W(t)=\sqrt{t^{2}+1}-1$ or $W(t)=\log\frac{e^{t}+e^{-t}}{2}$ with $(u,v)\in[-B,B]\times\mathcal{Y}$ and $\mathcal{Y}$ a compact subset of $\mathbb{R}$.
  • Cauchy loss [9]: $\ell(u,v)=\log\left(1+\frac{(u-v)^{2}}{2}\right)$, where $(u,v)\in[-B,B]\times\mathcal{Y}$, $B=\frac{1}{2}$, and $\mathcal{Y}=\left[-\frac{1}{2},\frac{1}{2}\right]$.
  • Welsch loss [27]: $\ell(u,v)=1-\exp\left(-\frac{(u-v)^{2}}{2}\right)$, where $(u,v)\in[-B,B]\times\mathcal{Y}$, $B=\frac{1}{3}$, and $\mathcal{Y}=\left[-\frac{1}{3},\frac{1}{3}\right]$.
Among these, the third loss function is the standard choice for Poisson regression. Notably, both the Cauchy and Welsch losses are globally non-convex, and the latter has attracted considerable attention in the image processing community [6].
2.2 Spherical Radial Basis Functions
[…]
$\mathcal{H}_{k}^{d}:=\left\{P\in\mathcal{P}_{k}^{d}\ |\ \Delta P=0\right\},$
where $\Delta$ is Laplacian operator. According to Chapter 1.2 of [18], the space $\mathcal{H}_{k}^{d}$ is a reproducing kernel Hilbert space (RKHS) with kernel $K_{k}(x,x^{\prime})=Q_{k}^{d}(\langle x,x^{\prime}\rangle)$ for $d\geq 3$, where $Q_{k}^{d}$ denotes the generalized-Legendre polynomial and $\langle x,x^{\prime}\rangle$ is the standard inner product in $\mathbb{R}^{d}$. When $d=2$, $\mathcal{H}_{k}^{d}$ is also an RKHS with kernel function $K_{k}(x,x^{\prime})$ given in Chapter 1.6.1 of [18]. The generalized Legendre polynomials $Q_{k}^{d}(u)$ for $d\geq 3$ are defined by
$Q_{1}^{d}(u)$ $:=1$
$\frac{1}{\Omega_{d-1}}\int_{-1}^{1}Q_{k}^{d}(u)Q_{j}^{d}(u)(1-u^{2})^{\frac{d-3}{2}}du$ $:=\frac{\dim\mathcal{H}_{k}^{d}}{\Omega_{d-2}}\delta_{k,j},\ \ \forall\ k,j\geq 1.$
For the orthonormal basis $\{Y_{k,j}\}_{1\leq j\leq\dim\mathcal{H}_{k}^{d}}$ of the space $(\mathcal{H}_{k}^{d},\langle\cdot,\cdot\rangle_{\omega})$, we have $K_{k}(x,x^{\prime})=\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}Y_{k,j}(x)Y_{k,j}(x^{\prime})$. Another important property is that the spaces $\{(\mathcal{H}_{k}^{d},\langle\cdot,\cdot\rangle_{\omega})\}_{k\geq 0}$ are mutually orthogonal and form an orthogonal decomposition of both $\mathcal{L}^{2}\left(\mathbb{S}^{d-1}\right)$ and $\Pi_{k}^{d}$, where $\bigoplus$ denotes the direct sum of inner product spaces,
$\Pi_{k}^{d}=\bigoplus_{0\leq j\leq k}\mathcal{H}_{j}^{d}\quad\quad\text{and}\quad\quad\mathcal{L}^{2}\left(\mathbb{S}^{d-1}\right)=\bigoplus_{k\geq 0}\mathcal{H}_{k}^{d}.$
We now introduce a common class of SBFs, $Q(u):=\sum_{k=0}^{\infty}a_{k}Q_{k}^{d}(u),$ which induces the kernel function
$K(x,x^{\prime}):=\sum_{k=0}^{\infty}a_{k}Q_{k}^{d}(\langle x,x^{\prime}\rangle)=\sum_{k=0}^{\infty}a_{k}K_{k}(x,x^{\prime})=\sum_{k=0}^{\infty}a_{k}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}Y_{k,j}(x)Y_{k,j}(x^{\prime}).$ (2.3)
The coefficients $0<a_{k}\leq 1$ satisfy $l:=\lim_{k\to\infty}a_{k}\cdot\left(\dim\Pi_{k}^{d}\right)^{2s}\in(0,\infty)$ for some $s>\frac{1}{2}$, with $\left(\dim\Pi_{k}^{d}\right)^{2s}=\mathcal{O}(k^{2s(d-1)})$. For such a kernel $K(x,x^{\prime})$, we established in Proposition A.1 of Subsection A.1 that $K(x,x^{\prime})$ converges uniformly and is therefore continuous. Together with its easily verifiable symmetry and positive semi-definiteness, $K(x,x^{\prime})$ is a Mercer kernel [40], inducing the RKHS $\mathcal{H}_{K}$ given by
$\mathcal{H}_{K}=\left\{f=\sum_{k=0}^{\infty}\sum_{1\leq j\leq\dim\mathcal{H}_{k}^{d}}f_{k,j}Y_{k,j}\,\Bigg|\,\sum_{k=0}^{\infty}\sum_{1\leq j\leq\dim\mathcal{H}_{k}^{d}}\frac{\left(f_{k,j}\right)^{2}}{a_{k}}<\infty\right\}\$(2.4)
with inner product
$\left\langle f,g\right\rangle_{K}:=\sum_{k=0}^{\infty}\sum_{1\leq j\leq\dim\mathcal{H}_{k}^{d}}\frac{f_{k,j}\cdot g_{k,j}}{a_{k}}.$ (2.5)
The capacity parameter $s$ is used to characterize the complexity of the hypothesis space $\mathcal{H}_{K}$, and as $s$ increases, the space $\mathcal{H}_{K}$ becomes smaller. Under the new inner product $\langle\cdot,\cdot\rangle_{K}$, the spaces $\{\mathcal{H}_{k}^{d}\}_{k\geq 0}$ remain mutually orthogonal. Moreover, each $\mathcal{H}_{k}^{d}$ is an RKHS with kernel $a_{k}K_{k}(x,x^{\prime})$ under $\langle\cdot,\cdot\rangle_{K}$. For further details on $\mathcal{H}_{K}$, we refer the reader to our previous work [5]. Given an increasing sequence of non-negative integers $\{L_{n}\}_{n\geq 0}\subset{\mathbb{N}}$, we define an increasing family of finite-dimensional, nested function spaces $\{\mathcal{H}_{L_{n}}\}_{n\geq 0}\subset\mathcal{H}_{K}$ by $\mathcal{H}_{L_{n}}:=\bigoplus_{k=0}^{L_{n}}\mathcal{H}_{k}^{d}$, as described in Section 1. According to Theorem 12.20 of [58] and the orthogonality of $\{\mathcal{H}_{k}^{d}\}_{k\geq 0}$, the space $\left(\mathcal{H}_{L_{n}},\langle\cdot,\cdot\rangle_{K}\right)$ forms an RKHS with kernel $K_{L_{n}}^{T}(x,x^{\prime})$, which expands as
$K_{L_{n}}^{T}(x,x^{\prime})=\sum_{k=0}^{L_{n}}a_{k}K_{k}(x,x^{\prime})=\sum_{k=0}^{L_{n}}a_{k}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}Y_{k,j}(x)Y_{k,j}(x^{\prime}),$ (2.6)
with inner product $\left\langle f,g\right\rangle_{K}=\sum_{k=0}^{L_{n}}\sum_{1\leq j\leq\dim\mathcal{H}_{k}^{d}}\frac{f_{k,j}\cdot g_{k,j}}{a_{k}}$ for all $f,g\in\mathcal{H}_{L_{n}}$.
2.3 Truncated Kernel Stochastic Gradient Descent
[…]
$\mathcal{W}:=\left\{f\in\mathcal{H}_{K}\,|\ \|f\|_{K}\leq Q\right\}.$ (2.7)
Hence, for all $f\in\mathcal{W}$, we have $\|f\|_{\infty}\leq\kappa Q<B$.
Under Assumption 1, Assumption 2, and the reproducing property of $\mathcal{H}_{K}$, Lemma A.1 yields the following inequality for the Fréchet derivative [15]. For any $f\in\mathcal{W}$ and $h\in\mathcal{H}_{K}$, it holds that
$o(\|h\|_{K})$ $=\mathbb{E}\left[\ell(f(X)+h(X),Y)-\ell(f(X),Y)-\partial_{u}\ell(f(X),Y)h(X)\right]$ (2.8)
$=\mathcal{E}(f+h)-\mathcal{E}(f)-\left\langle\mathbb{E}\left[\partial_{u}\ell(f(X),Y)K(X,\cdot)\right],h\right\rangle_{K}.$
The Fréchet derivative of $\mathcal{E}(f)$ in $\mathcal{H}_{K}$ is $\nabla\mathcal{E}(f)\big|_{\mathcal{H}_{K}}=\mathbb{E}\left[\partial_{u}\ell(f(X),Y)K(X,\cdot)\right]$, for which $\widehat{\nabla\mathcal{E}(f)}\big|_{\mathcal{H}_{K}}=\partial_{u}\ell(f(X_{n}),Y_{n})K(X_{n},\cdot)$ serves as an unbiased estimator.
We choose an increasing sequence of non-negative integers $\{L_{n}\}_{n\geq 0}$, typically defined as $L_{n}=\min\left\{k\,\Big|\,\dim{\Pi_{k}^{d}}\geq n^{\theta}\right\}$ with $0<\theta<\frac{1}{2}$. At the $n$-th iteration, we project the unbiased estimator $\widehat{\nabla\mathcal{E}(f)}\big|_{\mathcal{H}_{K}}$ onto the hypothesis space $\mathcal{H}_{L_{n}}=\bigoplus_{k=0}^{L_{n}}\mathcal{H}_{k}^{d}$ (see (2.6) for more details), given by
$P_{\mathcal{H}_{L_{n}}}\left(\widehat{\nabla\mathcal{E}(f)}\big|_{\mathcal{H}_{K}}\right)=\partial_{u}\ell(f(X_{n}),Y_{n})K_{L_{n}}^{T}(X_{n},\cdot),$
where $P_{\mathcal{H}_{L_{n}}}$ denotes the projection operator from $\mathcal{H}_{K}$ onto $\mathcal{H}_{L_{n}}$, and this result is established in Lemma A.2. Lemma A.2 also shows that for any $f\in\mathcal{H}_{L_{n}}\cap\mathcal{W}$, $\partial_{u}\ell(f(X_{n}),Y_{n})K_{L_{n}}^{T}(X_{n},\cdot)$ is an unbiased estimator of the gradient of the population risk $\mathcal{E}(f)$ in $\mathcal{H}_{L_{n}}$. In the algorithm, by tuning the parameter $\theta$, which determines the dimensionality of the hypothesis space $\mathcal{H}_{L_{n}}$, we establish a regularization mechanism that adapts to the complexity of $f^{*}$. Specifically, a smaller $\theta$ helps prevent overfitting when $f^{*}$ exhibits strong regularity, whereas a larger $\theta$ mitigates underfitting under weak regularity. In addition, we introduce the projection operator $P_{\mathcal{W}}:\mathcal{H}_{K}\rightarrow\mathcal{W}$, which projects elements of $\mathcal{H}_{K}$ onto $\mathcal{W}$ to ensure that each iteration remains in $\mathcal{W}$. This projection step is also standard in classical finite-dimensional stochastic gradient descent algorithms [33, 30]. Using unbiased estimates of the derivatives, we recursively define a sequence of iterates $\hat{f}_{n}\in\mathcal{H}_{L_{n}}\cap\mathcal{W}$, starting from the initialization $\hat{f}_{0}=0$, and
$\hat{f}_{n}$= $P_{\mathcal{W}}\left(\hat{f}_{n-1}-\gamma_{n}\partial_{u}\ell(\hat{f}_{n-1}(X_{n}),Y_{n})K_{L_{n}}^{T}(X_{n},\cdot)\right)$ (2.9)
$=$ $P_{\mathcal{W}}\left(\hat{f}_{n-1}-\gamma_{n}\partial_{u}\ell(\hat{f}_{n-1}(X_{n}),Y_{n})\sum_{k=0}^{L_{n}}a_{k}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}Y_{k,j}(X_{n})Y_{k,j}\right)$
with step size $\gamma_{n}=\gamma_{0}n^{-t}$ for $t\in\left[\frac{1}{2},1\right)$ and $\gamma_{0}>0$. In Lemma A.4, we show that $P_{\mathcal{W}}(f)\in\mathcal{H}_{L_{n}}\cap\mathcal{W}$ for any $f\in\mathcal{H}_{L_{n}}$. By induction, since $\hat{f}_{n-1}\in\mathcal{H}_{L_{n-1}}$ and $K_{L_{n}}^{T}(X_{n},\cdot)\in\mathcal{H}_{L_{n}}$, it follows that $\hat{f}_{n}\in\mathcal{W}\cap\mathcal{H}_{L_{n}}$. In Lemma A.6, we provide an explicit expression for the projection operator $P_{\mathcal{W}}$ in the subspace $\mathcal{H}_{L_{n}}$. For $f=\sum_{k=0}^{L_{n}}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}f_{k,j}Y_{k,j}\in\mathcal{H}_{L_{n}}$, we have
$P_{\mathcal{W}}(f)=\left\{\begin{aligned} &\frac{Q}{\|f\|_{K}}f=\frac{Q}{\left(\sum_{k=0}^{L_{n}}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}a_{k}^{-1}f_{k,j}^{2}\right)^{\frac{1}{2}}}f,\ \ \text{if}\ \|f\|_{K}>Q,\\ &f,\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\ \,\ \ \text{if}\ \|f\|_{K}\leq Q.\end{aligned}\right$. (2.10)
In addition to outputting the last iterate $\hat{f}_{n}$, T-kernel SGD also adopts a more robust $\alpha$-suffix averaging scheme. Specifically, for a fixed averaging parameter $\alpha\in(0,1)$, we define
$\bar{f}_{\alpha n}:=\frac{1}{\alpha n}\left(\hat{f}_{(1-\alpha)n}+\dots+\hat{f}_{n-2}+\hat{f}_{n-1}\right).$
Note that $\hat{f}_{n-1}\in\mathcal{H}_{L_{n-1}}$, we denote $\hat{f}_{n-1}=\sum_{k=0}^{L_{n}}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}f_{k,j}^{(n-1)}Y_{k,j}$ (with $f_{L_{n},j}^{(n-1)}=0$) and define $\hat{g}_{n}:=\hat{f}_{n-1}-\gamma_{n}\partial_{u}\ell(\hat{f}_{n-1}(X_{n}),Y_{n})K_{L_{n}}^{T}(X_{n},\cdot)$. In practice, the update of $\hat{g}_{n}$ is performed directly on the coefficients of its expansion, i.e.,
$\hat{g}_{n}=\sum_{k=0}^{L_{n}}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}g_{k,j}^{(n)}Y_{k,j}:=\sum_{k=0}^{L_{n}}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}\left(f_{k,j}^{(n-1)}-\gamma_{n}\partial_{u}\ell(\hat{f}_{n-1}(X_{n}),Y_{n})a_{k}Y_{k,j}(X_{n})\right)Y_{k,j}.$
From equation (2.10), the projection operation on $\hat{g}_{n}$, i.e., $\hat{f}_{n}=P_{\mathcal{W}}(\hat{g}_{n})$, essentially only involves operations on the coefficients of the expansion of $\hat{g}_{n}$. In the recursion of the T-Kernel SGD (2.9), aside from computing the function value $\hat{f}_{n}(X_{n})=\sum_{k=0}^{L_{n}}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}f_{k,j}^{(n)}Y_{k,j}(X_{n})$, all other operations are performed on the coefficients of the basis functions $\{Y_{k,j}\}$. The explicit forms of the basis functions $\{Y_{k,j}\}$, as well as the normalization constants and related details, are provided in subsubsection A.1.1. Therefore, we can directly present the T-Kernel SGD, which works with the coefficients of the expansion, in Algorithm 1.
Algorithm 1 Truncated Kernel Stochastic Gradient Descent
set: $s>\frac{1}{2},\ \gamma_{0}>0,\ \frac{1}{2}\leq t<1$, $0<\theta<\frac{1}{2}$, and $L_{0}=0$.
initialize: $\hat{f}_{0}=0,\ K^{T}_{L_{0}}(x,\cdot)=a_{0}K_{0}(x,\cdot)=a_{0}Y_{0,1}(x)Y_{0,1}$.
for $n=1,2,3,\dots$ do
  Collect sample $(X_{n},Y_{n})$, calculate $\gamma_{n}=\gamma_{0}n^{-t}$ and $L_{n}$.
  Update $\hat{g}_{n}$ :
$\hat{g}_{n}=$ $\hat{f}_{n-1}-\gamma_{n}\partial_{u}\ell(\hat{f}_{n-1}(X_{n}),Y_{n})K_{L_{n}}^{T}(X_{n},\cdot)$
$=$ $\sum_{k=0}^{L_{n}}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}g_{k,j}^{(n)}Y_{k,j}:=\sum_{k=0}^{L_{n}}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}\left(f_{k,j}^{(n-1)}-\gamma_{n}\partial_{u}\ell(\hat{f}_{n-1}(X_{n}),Y_{n})a_{k}Y_{k,j}(X_{n})\right)Y_{k,j}.$
  Calculate $\|\hat{g}_{n}\|_{K}^{2}=\sum_{k=0}^{L_{n}}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}a_{k}^{-1}\left(g_{k,j}^{(n)}\right)^{2}$ and update $\hat{f}_{n}$:
$\hat{f}_{n}=P_{\mathcal{W}}(\hat{g}_{n})=\left\{\begin{aligned} &\frac{Q}{\|\hat{g}_{n}\|_{K}}\cdot\sum_{k=0}^{L_{n}}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}g_{k,j}^{(n)}Y_{k,j},\ \ \text{if}\ \|\hat{g}_{n}\|_{K}>Q,\\ &\hat{g}_{n},\ \ \text{if}\ \|\hat{g}_{n}\|_{K}\leq Q.\end{aligned}\right$.
  $n\leftarrow n+1$
end for
return $\hat{f}_{n},\bar{f}_{\alpha n}=\frac{1}{\alpha n}\sum_{i=(1-\alpha)n}^{n-1}\hat{f}_{i}$
In Algorithm 1, the computational cost of each update is mainly attributed to evaluating $\hat{f}_{n-1}(X_{n})$, updating $\hat{g}_{n}$, $\hat{f}_{n}$, and computing $\|\hat{g}_{n}\|_{K}$. The latter three operations require comparable computational time $\mathcal{O}\left(\sum_{k=0}^{L_{n}}\dim\mathcal{H}_{k}^{d}\right)=\mathcal{O}\left(\dim\Pi_{L_{n}}^{d}\right)$. The former, however, requires computing the basis functions $\{Y_{k,j}(X_{n})\}$. As shown in subsubsection A.1.1, the evaluation of each basis function $\{Y_{k,j}(X_{n})\}$ for $0\leq k\leq L_{n}$ can be performed in at most $\mathcal{O}(dL_{n})$ time, which implies that the evaluation of $\hat{f}_{n-1}(X_{n})$ takes at most $\mathcal{O}(dL_{n}\dim\Pi_{L_{n}}^{d})$ time. Consequently, the total computational cost of T-Kernel SGD for processing $n$ samples is $\mathcal{O}(dnL_{n}\dim\Pi_{L_{n}}^{d})$. In terms of storage complexity, T-kernel SGD only requires maintaining the coefficients of $\hat{f}_{n}$ and $\hat{g}_{n}$, together with intermediate quantities represented in the coefficients of the basis functions $\{Y_{k,j}\}_{0\leq k\leq L_{n},1\leq j\leq\dim\mathcal{H}_{k}^{d}}$. Consequently, the memory consumption of the algorithm is $\mathcal{O}(\dim\Pi_{L_{n}}^{d})$. A more in-depth analysis of both computational and storage complexities is provided in Subsection 3.1.
[…]
$\|f\circ F\|_{\rho_{X}}^{2}=\int_{\Omega}|f\circ F(X)|^{2}d\rho_{X}\leq\|f\|_{\infty}^{2}\leq\kappa^{2}\|f\|_{K}^{2}.$
We still consider the population risk minimization problem
$f^{*}:=\arg\min_{f\in\mathcal{W}}\mathcal{E}(f)=\arg\min_{f\in\mathcal{W}}\mathbb{E}_{\rho}\left[\ell\left(f\circ F\left(X\right),Y\right)\right].$
Through the mapping $F$, we establish a generalized T-kernel SGD algorithm for non-spherical samples, starting from the initialization $\hat{f}_{0}=0$, and
$\hat{f}_{n}$= $P_{\mathcal{W}}\left(\hat{f}_{n-1}-\gamma_{n}\partial_{u}\ell(\hat{f}_{n-1}\circ F(X_{n}),Y_{n})K_{L_{n}}^{T}(F(X_{n}),\cdot)\right),$ (2.11)
we adopt the same parameter settings as in the original T-kernel SGD, namely $L_{n}=\min\newline \left\{k\,\Big|\,\dim{\Pi_{k}^{d}}\geq n^{\theta}\right\}$ with $0<\theta<\frac{1}{2}$ and $\gamma_{n}=\gamma_{0}n^{-t}$ for $t\in\left[\frac{1}{2},1\right)$. In addition, we employ the $\alpha$-suffix averaging scheme $\bar{f}_{\alpha n}:=\frac{1}{\alpha n}\sum_{i=(1-\alpha)n}^{n-1}\hat{f}_{i}$ as the output of the estimator.
3 Theoretical Results
This section focuses on establishing the optimal generalization guarantees of the generalized T-kernel SGD algorithm (2.11). Our analysis builds on the notions introduced at the end of Subsection 2.3, including the mapping $F$ and the unknown distribution $\rho$. Before stating the assumptions on $\rho$ and the minimizer $f^{*}$, we introduce the covariance operator
$L_{\omega,K}:\mathcal{L}^{2}\left(\mathbb{S}^{d-1}\right)$ $\to\mathcal{L}^{2}\left(\mathbb{S}^{d-1}\right)$
[…]
By the definition of $K(x,x^{\prime})$ in (2.3) and the fact that $\{Y_{k,j}\}_{0\leq k,\;1\leq j\leq\dim\mathcal{H}_{k}^{d}}$ forms an orthonormal basis of $\mathcal{L}^{2}\left(\mathbb{S}^{d-1}\right)$, the covariance operator $L_{\omega,K}$ admits an orthonormal eigensystem $\left\{\left(a_{k},Y_{k,j}\right)\right\}_{0\leq k,1\leq j\leq\dim\mathcal{H}_{k}^{d}}$. Next, for any $r\geq\frac{1}{2}$, the $r$-th power of $L_{\omega,K}$, denoted by $L_{\omega,K}^{r}$, is defined by
$L_{\omega,K}^{r}:\mathcal{L}^{2}\left(\mathbb{S}^{d-1}\right)$ $\to\mathcal{L}^{2}\left(\mathbb{S}^{d-1}\right)$
$\sum_{k=0}^{\infty}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}f_{k,j}Y_{k,j}$ $\to\sum_{k=0}^{\infty}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}a_{k}^{r}f_{k,j}Y_{k,j}.$
Assumption 4.
The samples ${(X_{i},Y_{i})}_{i\in{\mathbb{N}}+}\in\Omega\times\mathcal{Y}$ are independently and identically distributed (i.i.d.) according to the Borel probability distribution $\rho$.
Assumption 5.
(Regularity condition) The minimizer $f^{*}$, defined as
$f^{*}:=\arg\min_{f\in\mathcal{W}}\mathcal{E}(f)=\arg\min_{f\in\mathcal{W}}\mathbb{E}_{\rho}\left[\ell\left(f\circ F\left(X\right),Y\right)\right],$
satisfies $f^{*}=L_{\omega,K}^{r}\left(g^{*}\right)$ for some $r\geq\frac{1}{2}$ and $g^{*}\in\mathcal{L}^{2}\left(\mathbb{S}^{d-1}\right)$. Moreover, $f^{*}$ fulfills one of the following conditions:
  • (a).
    $f^{*}$ lies in the interior of $\mathcal{W}$, i.e., $\|f^{*}\|_{K}<Q$.
  • (b).
    There exists a constant $L>0$ such that, for every $f\in\mathcal{W}$,
    $\mathcal{E}(f)-\mathcal{E}(f^{*})\leq\frac{L}{2}\left\|f\circ F-f^{*}\circ F\right\|_{\rho_{X}}^{2}.$ (3.1)
Assumption 6.
The marginal distribution $\rho_{X}$ is absolutely continuous with respect to the Lebesgue measure $\lambda$ on $\Omega$, with the Radon–Nikodym derivative $\frac{d\rho_{X}}{d\lambda}$. Moreover, there exist constants $0<b_{\rho}^{\prime}<B_{\rho}^{\prime}$ such that
[…]
The regularity condition stated in Assumption 5 is a key assumption on the smoothness of the minimizer $f^{*}$ and is standard in the online-learning literature [53, 63, 56, 19, 23]. In fact, a larger $r$ means that the expansion coefficients $\{\langle f^{*},Y_{k,j}\rangle\}$ of $f^{*}$ decay more rapidly, indicating stronger regularity of $f^{*}$. According to Theorem 4 in [17], if $r\geq\tfrac{1}{2}$ then $L_{\omega,K}^{r}(\mathcal{L}^{2}\left(\mathbb{S}^{d-1}\right))\subset\mathcal{H}_{K}$, and more generally $L_{\omega,K}^{r_{1}}(\mathcal{L}^{2}\left(\mathbb{S}^{d-1}\right))\subset L_{\omega,K}^{r_{2}}(\mathcal{L}^{2}\left(\mathbb{S}^{d-1}\right))$ for all $r_{1}\geq r_{2}$. In the finite-dimensional setting, condition (b) of Assumption 5 is a special case of the descent lemma for $L$-smooth functions [41, 8]. By analogy, in our analysis we combine condition (a) of Assumption 5 with the $L$-smoothness property and, invoking Lemma A.3, establish the inequality stated in (b). Therefore, we do not distinguish between the Lipschitz constant $L$ in Assumption 2 and the constant $L$ in (b) of Assumption 5.
Compared with the assumptions on the unknown distribution $\rho$ in previous work on nonparametric regression [54, 13, 19, 23], Assumption 6 is more direct. In particular, Assumption 6 plays a key role in establishing the equivalence between the two norms $\|\cdot\|_{\rho_{X}}$ and $\|\cdot\|_{\omega}$. As shown in Lemma A.7, there exist constants $0<b_{\rho}<B_{\rho}$ such that
$b_{\rho}\Omega_{d-1}\|f\|_{\omega}^{2}\leq\|f\circ F\|_{\rho_{X}}^{2}\leq B_{\rho}\Omega_{d-1}\|f\|_{\omega}^{2},\quad\forall f\in\mathcal{H}_{K}.$ (3.3)
This inequality is crucial for deriving one of the central tools of this paper—the asymptotic equivalence between the RKHS norm $\|\cdot\|_{K}$ and the distribution-dependent norm $\|\cdot\|_{\rho_{X}}$.
3.1 Optimal Rates for Excess Risk
Our first main result provides rate-optimal convergence guarantees for the expected excess risk, $\mathbb{E}\left[\mathcal{E}(\hat{f}_{n})-\mathcal{E}(f^{*})\right]$, where $\hat{f}_{n}$ denotes the estimator produced by T-kernel SGD under general loss functions.
Theorem 1.
Assume that Assumption 1 (with $M>0$), Assumption 2 (with $L>0$), Assumption 3 (with $\mu>0$), Assumption 4, Assumption 5 (with $r\geq\frac{1}{2}$), and Assumption 6 (with $0<b_{\rho}<B_{\rho}$ in (3.3)) hold. Let $\theta=\frac{1}{2s(2r+1)}$ and choose the step size $\gamma_{n}=\gamma_{0}n^{-\frac{2r}{2r+1}}\log(n+1)$, where $\gamma_{0}=c\frac{A_{1}4(2d)^{2s}}{A^{2}_{2}b_{\rho}\mu\Omega_{d-1}}$ for some constant $c\in\left[\frac{1}{\log 2},\frac{2}{\log 3}\right]$. Then, for any $\alpha\in(0,1)$, the following bounds hold:
$\mathbb{E}\left[\mathcal{E}\left(\hat{f}_{n}\right)-\mathcal{E}\left(f^{*}\right)\right]$ $\leq\mathcal{O}\left(n^{-\frac{2r}{2r+1}}\left(\log(n+1)\right)^{2}\right),$
$\mathbb{E}\left[\mathcal{E}\left(\bar{f}_{\alpha n}\right)-\mathcal{E}\left(f^{*}\right)\right]$ $\leq\mathcal{O}\left(n^{-\frac{2r}{2r+1}}\log(n+1)\right),$
where $\hat{f}_{n}$ denotes the last iterate in (2.11) and $\bar{f}_{\alpha n}=\frac{1}{\alpha n}\sum_{i=(1-\alpha)n}^{n-1}\hat{f}_{i}$ is the $\alpha$-suffix average. Here, $0<A_{2}\leq 1\leq A_{1}$ denote the upper and lower bounds, of $a_{k}\cdot\left(\dim\Pi_{k}^{d}\right)^{2s}$, respectively, i.e.,
$A_{2}\left(\dim\Pi_{k}^{d}\right)^{-2s}\leq a_{k}\leq A_{1}\left(\dim\Pi_{k}^{d}\right)^{-2s},\ \ \forall k\in{\mathbb{N}}.$
In online nonparametric regression, most studies on minimax optimality have focused almost exclusively on the least-squares loss, whereas investigations of general loss function, particularly non-convex ones, remain limited. Nevertheless, classical kernel SGD typically suffers from the saturation phenomenon, where the convergence rate ceases to improve once the regularity of the minimizer $f^{*}$ exceeds a certain threshold. For unregularized kernel SGD, [63] established convergence rates of $\mathcal{O}\left(n^{-\frac{2r}{2r+1}}\log n\right)$ for the regularity parameter $r\in(0,\frac{1}{2}]$, while [24] obtained optimal rates $\mathcal{O}\left(n^{-\frac{2r}{2r+1}}\right)$ using the capacity parameter $s$, valid for $r\in\left[\frac{1}{2},1-\frac{1}{4s}\right]$. By employing Polyak averaging, [19] enhanced the robustness of the estimator and established optimal convergence rates $\mathcal{O}\left(n^{-\frac{4sr}{4sr+1}}\right)$, which depend on the capacity parameter $s$, for $r\in\left[\frac{1}{2}-\frac{1}{4s},1-\frac{1}{4s}\right]$. Incorporating an additional regularization scheme into kernel SGD helps alleviate saturation. In particular, [56] analyzed regularized kernel SGD and obtained the optimal rates $\mathcal{O}\left(n^{-\frac{2r}{2r+1}}\left(\log\frac{2}{\alpha}\right)^{4}\right)$ with probability at least $1-\alpha$ for $r\in\left[\frac{1}{2},1\right]$. In contrast to previous analyses, which experience saturation when the regularity parameter $r>1$, our algorithm, when specialized to the least-squares case, effectively overcomes this phenomenon. For general loss functions, however, the nonlinear structure of the Hessian introduces substantial challenges in analyzing the convergence of kernel SGD. In online learning, classical SGD analysis yields only the slow rate $\mathcal{O}(n^{-\frac{1}{2}})$, corresponding to saturation at $r=\frac{1}{2}$. Leveraging stronger regularity conditions ($r>\frac{1}{2}$) to accelerate kernel SGD has remained an open problem. Theorem 1 shows that, under suitable regularity assumptions, T-kernel SGD attains fast rates and—to the best of our knowledge—provides the first saturation-free guarantees for online learning with general losses.
In T-kernel SGD, choosing an appropriate size for the hypothesis space $\mathcal{H}_{L_{n}}$ is crucial to achieving optimal rates. When the minimizer $f^{*}$ exhibits higher smoothness, i.e., when the regularity parameter $r$ is larger, a smaller $\theta$ should be selected to reduce variance; conversely, for a less smooth $f^{*}$, a larger $\theta$ is preferable to control bias. Accordingly, in Theorem 1 we set $\theta=\tfrac{1}{2s(2r+1)}$, which effectively balances bias and variance. In contrast, the regularization strategies commonly used in classical kernel SGD, such as approximating the regularization path or tuning the step size—affect the complexity of the hypothesis space only indirectly and to a limited extent. As a result, when the minimizer $f^{*}$ has regularity $r>1$, these methods fail, leading to the saturation phenomenon. Moreover, the finite-dimensional structure of $\mathcal{H}_{L_{n}}$ is essential for the convergence analysis. Building on the norm equivalence between $\|\cdot\|_{\rho_{X}}$ and $\|\cdot\|_{\omega}$ shown in (3.3), we further establish the asymptotic equivalence between $\|\cdot\|_{\rho_{X}}$ and $\|\cdot\|_{K}$ (see Lemma A.11),
$\frac{A^{2}_{2}}{A_{1}}\frac{b_{\rho}\Omega_{d-1}}{(2d)^{2s}}\,n^{-2\theta s}\|f\|_{K}^{2}\leq\|f\circ F\|_{\rho_{X}}^{2}\leq\kappa^{2}\|f\|_{K}^{2},\ \ \forall f\in\mathcal{H}_{L_{n}}.$
Furthermore, the asymptotic equivalence above serves as an inequality-based characterization of the covariance operator $L_{\omega,K}$, effectively capturing the decay rate of its eigenvalues. By combining optimization techniques with this inequality and the inequality-based characterization of the regularity of the minimizer $f^{*}$ in Lemma A.12, we present the proof of Theorem 1 in Subsection A.3. Applying the local strong convexity of losses, we establish the following result in Subsection A.4.
Proposition 1.
Suppose the assumptions in Theorem 1 hold. Choose $\theta=\frac{1}{2s(2r+1)}$ and set the step size $\gamma_{n}=\gamma_{0}n^{-\frac{2r}{2r+1}}\log(n+1)$, where $\gamma_{0}=c\frac{A_{1}4(2d)^{2s}}{A^{2}_{2}b_{\rho}\mu\Omega_{d-1}}$ for some constant $c\in\left[\frac{1}{\log(2)},\frac{2}{\log(3)}\right]$. Then, for any $\alpha\in(0,1)$, we have
$\mathbb{E}\left[\left\|\hat{f}_{n}\circ F-f^{*}\circ F\right\|_{\rho_{X}}^{2}\right]$ $\leq\mathcal{O}\left(n^{-\frac{2r}{2r+1}}\left(\log(n+1)\right)^{2}\right)$
$\mathbb{E}\left[\left\|\bar{f}_{\alpha n}\circ F-f^{*}\circ F\right\|_{\rho_{X}}^{2}\right]$ $\leq\mathcal{O}\left(n^{-\frac{2r,}{2r+1}}\log(n+1)\right).$
In Proposition 1, we show that convergence of the excess risk is equivalent to convergence in the $\|\cdot\|_{\rho_{X}}$ norm. Compared with the convergence in the RKHS discussed in the next subsection, this result can be interpreted as weak convergence.
We now turn to a more in-depth analysis of the computational and storage complexities of T-kernel SGD, and in particular demonstrate the optimality of the memory. Applying Lemma 2 and Lemma 4 in [5], we obtain $\dim\Pi_{L_{n}}^{d}\leq dn^{\theta}$ and $L_{n}\leq((d-1)!\dim\Pi_{L_{n}}^{d})^{1/(d-1)}$. Combining these bounds with the computational and storage complexity of T-kernel SGD derived in Subsection 2.3, we conclude that processing $n$ samples requires $\mathcal{O}(d^{\frac{3d-2}{d-1}}n^{1+\frac{d}{d-1}\theta})$ computational time and $\mathcal{O}(dn^{\theta})$ memory. In complexity analysis, both the computational and storage upper bounds grow rapidly with spherical dimension $d$. However, in practice, the orthonormal basis representations of low-order polynomial spaces often admit simplifications in high-dimensional settings (see subsubsection A.1.1 for details). Consequently, the actual complexity of the algorithm does not increase as drastically with $d$ as the theoretical bounds might suggest, a fact further supported by the high-dimensional experiments presented in Subsection 4.3. In Theorem 1, by choosing $\theta=\frac{1}{2s(2r+1)}$, the computational time is $\mathcal{O}(d^{\frac{3d-2}{d-1}}n^{1+\frac{d}{d-1}\frac{1}{2s(2r+1)}})$, while the memory requirement is $\mathcal{O}(dn^{\frac{1}{2s(2r+1)}})$. This is significantly lower than the computational cost $\mathcal{O}(n^{2})$ and the memory requirement $\mathcal{O}(n)$ of classical kernel SGD. To the best of our knowledge, T-kernel SGD achieves the highest computational efficiency among algorithms applicable to general losses, attaining the minimax optimal rates with the lowest time and memory complexities.
Since computers cannot store real numbers with infinite precision and typically represent data using finite binary sequences, additional errors may arise. To mitigate the impact of such errors on the optimality of the algorithm, one may increase the precision during the recursion. For instance, by employing binary sequences of length $2\log(n)$, a precision of order $\mathcal{O}(\frac{1}{n^{2}})$ can be achieved. Recently, [64] introduced a modified stochastic gradient descent algorithm that stores coefficients with a precision that increases with the sample size $n$. This method requires only an additional $\log(n)$ factor in the original storage complexity and achieves the theoretically optimal convergence rate. Thus, by making a simple modification to Algorithm 1, we can design an algorithm that gradually increases the coefficient precision, while requiring only an additional $\log(n)$ memory. Consequently, the storage complexity of the modified algorithm becomes $\mathcal{O}(dn^{\tfrac{1}{2s(2r+1)}}\log(n))$. In practice, however, the 64-bit double-precision floating-point representation (as used in Python) is typically sufficient for the implementation of the T-kernel SGD, and we therefore provide only a brief explanation here.
In the following, we investigate the optimality of the storage complexity. The relevant definitions and concepts employed in the discussion of the lower bound on storage complexity are adapted from Section 6.3 of [64]. We now adopt a description analogous to a (probabilistic) Turing machine to formally define the general estimator. An estimator can be viewed as a mapping $G_{n}$ from the sample space $\{(X_{i},Y_{i})\}_{1\leq i\leq n}\subset\left(\Omega\times\mathcal{Y}\right)^{n}$ to the function space $f_{n}\in\mathcal{W}$. Any estimator implementable on a computer necessarily involves an encoding–decoding procedure: the encoder $E_{n}$ maps the samples $\{(X_{i},Y_{i})\}_{1\leq i\leq n}$ to a binary sequence $b_{n}$, which is stored in memory, and the decoder $D_{n}$ translates the stored $b_{n}$ into the output function $\hat{f}_{n}$. In general, as the sample size increases, the estimator produces more accurate outputs, which in turn leads to an increase in the length of the binary sequence $b_{n}$. More specifically, we introduce the following definition of the general estimator.
Definition 1.
For $l_{n}\in{\mathbb{N}}_{+}$, we define an $l_{n}$-sized estimator $G_{n}=D_{n}\circ E_{n}:(\Omega\times\mathcal{Y})^{n}\to\mathcal{W}$, that is, the composition of the encoder $E_{n}$ and the decoder $D_{n}$.
  • (a).
    For $n\in{\mathbb{N}}_{+}$, one may consider an encoding map $E_{n}:(\Omega\times\mathcal{Y})^{n}\to\{0,1\}^{l_{n}}$, which can be randomized or deterministic.
  • (b).
    The decoder $D_{n}:\{0,1\}^{l_{n}}\to\mathcal{W}$ is a known, deterministic map that maps a binary sequence of length $l_{n}$ to a function in $\mathcal{W}$.
By combining the above definitions, one can derive a lower bound on the storage complexity while achieving the minimax rate.
Lemma 1.
Consider a positive integer sequence $\{l_{n}\}$ such that $l_{n}=o\left(n^{\frac{1}{2s(2r+1)}}\right)$ with $s>\frac{1}{2},\ r\geq\frac{1}{2}$, and let $G(l_{n})$ denote the collection of all $l_{n}$-sized estimators, one has
$\lim_{n\to\infty}\inf_{G_{n}\in G(l_{n})}\sup_{f^{*}\in\mathcal{W}\cap L_{\omega,K}^{r}\left(\mathcal{L}^{2}\left(\mathbb{S}^{d-1}\right)\right)}\mathbb{E}\left[n^{\frac{2r}{2r+1}}\|G_{n}\left(\{(X_{i},Y_{i})\}_{1\leq i\leq n}\right)-f^{*}\|_{\omega}^{2}\right]=\infty.$
The proof of Lemma 1 is provided in Subsection A.5. Lemma 1 implies that no estimator can achieve the optimal convergence rate while using memory of order $o\left(n^{\frac{1}{2s(2r+1)}}\right)$; that is, $\mathcal{O}\left(n^{\frac{1}{2s(2r+1)}}\right)$ constitutes a lower bound on the storage complexity. Consequently, taking into account the errors introduced by finite-precision memory, T-kernel SGD attains the optimal storage complexity up to a logarithmic factor $\log(n)$, which is substantially lower than the $\mathcal{O}(n)$ memory required by classical kernel SGD.
3.2 Optimal Rates for Strong Convergence
Our second main result, concerning convergence in the RKHS, often referred to as strong convergence, is presented below.
Theorem 2.
If the assumptions in Theorem 1 hold, choose $\theta=\frac{1}{2s(2r+1)}$ and set the step size $\gamma_{n}=\gamma_{0}n^{-\frac{2r}{2r+1}}\log(n+1)$, where $\gamma_{0}=c\frac{A_{1}4(2d)^{2s}}{A^{2}_{2}b_{\rho}\mu\Omega_{d-1}}$ for some constant $c\in\left[\frac{1}{\log(2)},\frac{2}{\log(3)}\right]$. Then, we have
$\mathbb{E}\left[\left\|\hat{f}_{n}-f^{*}\right\|_{K}^{2}\right]$
$\leq$ $\left(2Q^{2}+3A_{1}^{2r-1}\|g^{*}\|_{\omega}^{2}\right)(n+1)^{-\frac{2r-1}{2r+1}}+(4r+2)P^{2}(\log(n+1))^{2}(n+1)^{-\frac{2r-1}{2r+1}}.$
where $P^{2}$ is a constant given by
$P^{2}=\gamma_{0}^{2}\left[\left(\left(\frac{\mu}{2}+\frac{8L^{2}}{\mu}\right)\frac{L}{\mu}+L\right)B_{\rho}\Omega_{d-1}A_{1}^{2r}\|g^{*}\|_{\omega}^{2}\frac{1}{\gamma_{0}\log(2)}+M^{2}\kappa^{2}\right].$
By the inequality $\|f\|_{\infty}\leq\kappa\|f\|_{K}$, strong convergence guarantees uniform convergence. Moreover, as shown in [54], if the kernel $K\in C^{2k}(\Omega\times\Omega)$, then strong convergence implies convergence in the $C^{k}(\Omega)$ norm, where $C^{k}(\Omega)$ denotes the space of $k$-order continuously differentiable functions on $\Omega$, equipped with the norm $\|f\|_{C^{k}(\Omega)}=\sum_{|\alpha|\leq k}\|D^{\alpha}f\|_{\infty}$. Accordingly, strong convergence typically implies that the estimator approximates both the minimizer itself and its higher-order derivatives. Previous work has established strong convergence in various settings, including least-squares regression [63, 56, 24] and robust regression [23]. However, the above analyses are based on the classical kernel SGD algorithm, which requires handling all sample pairs $\{(X_{i},X_{j})\}_{1\leq i<j\leq n}$, leading to computational complexity $\mathcal{O}(n^{2})$ and storage complexity $\mathcal{O}(n)$. Such excessive costs severely limit its applicability to large-scale problems. Moreover, existing large-scale kernel methods [48, 1] have focused primarily on convergence in excess risk, leaving the development of efficient algorithms that achieve optimal strong convergence rates largely unexplored. In contrast, our work establishes T-kernel SGD, which is both computationally and memory efficient, and achieves capacity-dependent optimal rates (see, e.g., [10]) for strong convergence up to logarithmic factors.
4 Numerical Experiments
In Subsection 4.1 and Subsection 4.2, we demonstrate the theoretical analysis on two- and three-dimensional spheres, respectively, and conduct comparative experiments with the classical kernel SGD algorithm. In Subsection 4.3, we further evaluate the performance of the T-kernel SGD on the real high-dimensional MNIST dataset.
4.1 Robust Regression on the Circle
In this section, we validate the theoretical results presented in Section 3 by selecting optimal functions $f^{*}$ that satisfy different regularity conditions. In the experiments, we consider three classical loss functions commonly employed in robust regression: Cauchy, Huber, and Welsch losses. The experimental results demonstrate that T-kernel SGD effectively overcomes the saturation issue, attaining minimax rates that surpass the rate $\mathcal{O}(n^{-1/2})$. Moreover, compared with classical kernel SGD, it offers substantial improvements in computational efficiency.
In this subsection, we consider the model $Y=f^{*}(X)+\varepsilon,$ where $X$ is uniformly distributed on $\mathbb{S}^{1}$, and the noise term $\varepsilon$ is also uniformly distributed. Let $x=(\cos\theta,\sin\theta)$, $x^{\prime}=(\cos\varphi,\sin\varphi)\in\mathbb{S}^{1}$, and consider the following kernel for T-kernel SGD:
$K(x,x^{\prime})$ $=K_{0}(x,x^{\prime})+\sum_{k=1}^{\infty}\frac{1}{(2k)^{2s}}K_{k}(x,x^{\prime})\overset{\text{(i)}}{=}1+\sum_{k=1}^{\infty}\frac{2}{(2k)^{2s}}\cos(k(\theta-\varphi))$ (4.1)
$\overset{\text{(ii)}}{=}1+\frac{\sqrt{2}(-1)^{s+1}\pi^{2s}}{2(2s)!}B_{2s}(\{\tfrac{\theta-\varphi}{2\pi}\}),$
where $\{\theta\}$ denotes the fractional part of $\theta$, and $B_{2s}$ denotes the $2s$-th Bernoulli polynomial for $s\in\mathbb{N}$. For the details of equations (i) and (ii), see [18, 19]. According to Section 1.6.1 of [18], $\dim\mathcal{H}_{k}^{2}=2$ for $k\geq 1$. Consequently, the kernel $K_{k}(x,x^{\prime})$ on the two-dimensional sphere can be written as $K_{k}(x,x^{\prime})=Y_{k}^{1}(x)Y_{k}^{1}(x^{\prime})+Y_{k}^{2}(x)Y_{k}^{2}(x^{\prime}),$ and the orthonormal basis functions $Y_{k}^{1}$ and $Y_{k}^{2}$ admit simple explicit expressions, corresponding to the first- and second-kind Chebyshev polynomials, respectively. Therefore, each $\hat{f}_{n}$ can be explicitly represented as a truncated series $\hat{f}_{n}=\sum_{k=0}^{L_{n}}f_{k}Y_{k}^{1}+f^{\prime}_{k}Y_{k}^{2},$ and, when combined with iteration (2.9), only the coefficients of the truncated series need to be updated. Simultaneously, we choose $\mathcal{W}$ to be the closed unit ball of radius $Q=1$. For kernel SGD, we adopt a recursion similar to [32, 53, 63, 24], with the step size $\gamma_{n}=\gamma_{0}n^{-t}$:
$g_{n}=g_{n-1}-\gamma_{n}\,\partial_{u}\ell(g_{n-1}(X_{n}),Y_{n})K(X_{n},\cdot).$
In the comparative experiments of kernel SGD, we consider three different kernels: the Bernoulli polynomial kernel $\frac{\pi^{2}}{4}B_{2}$ and two widely used universal kernels, namely the Gaussian kernel and the Matérn-$\tfrac{5}{2}$ kernel. Let $r=\|x-x^{\prime}\|$, and the Gaussian and Matérn-$\tfrac{5}{2}$ kernels are given as follows:
$K_{Gaussian}(r)=\exp\left(\frac{r^{2}}{2}\right),\quad K_{Matern}^{5/2}(r)=\left(1+\sqrt{5}r+\frac{5r^{2}}{3}\right)\exp\left(-\sqrt{5}r\right).$
See Table 1 for the model setup.
  Example 1   Example 2
  $s$   1   1
  $r$   $\frac{7}{4}$   $\frac{3}{4}$
  optimal fitting $f^{*}$   $\frac{1}{2}B_{4}\left(\frac{\theta}{2\pi}\right)$   $\frac{1}{5}B_{2}\left(\frac{\theta}{2\pi}\right)$
  T-kernel SGD step size $\frac{\gamma_{n}}{\gamma_{0}}$   $n^{-7/9}$   $n^{-3/5}$
  kernel SGD step size $\frac{\gamma_{n}}{\gamma_{0}}$   $n^{-7/9}$   $n^{-3/5}$
  noise $\epsilon$   U[0.2,0.2]U[-0.2,0.2]   U[0.2,0.2]U[-0.2,0.2]
  Truncation level $L_{n}$   $n^{\frac{1}{9}}$   $n^{\frac{1}{5}}$
Table 1: Examples
Figure 1: The left figure illustrates the convergence of the error with respect to the sample size under three different losses, while the right figure shows the convergence of the error with respect to runtime. The black line indicates the minimax rate, with the slope $\frac{7}{9}$.
The comparative experimental results between kernel SGD and T-kernel SGD for Example 1 are presented in Figure 1. When the optimal function $f^{*}$ satisfies stronger regularity conditions ($r=\tfrac{7}{4}>1$), T-kernel SGD consistently achieves the theoretically optimal rate, and for non-convex losses (such as the Cauchy and Welsch losses), the algorithm still demonstrates strong performance. Moreover, it is noteworthy that kernel SGD exhibits clear saturation when using the Bernoulli polynomial kernel, with a convergence rate significantly slower than the minimax rate. Compared to kernel SGD, T-kernel SGD significantly improves computational efficiency. Owing to these gains in computational complexity, it substantially reduces training time while achieving superior convergence performance in a much shorter runtime.
Figure 2: The left figure illustrates the convergence of the error with respect to the sample size under three different losses, while the right figure shows the convergence of the error with respect to runtime. The black line indicates the minimax rate, with the slope $\frac{3}{5}$.
The experimental results for Example 2 are shown in Figure 2, demonstrating the convergence of the algorithm when $f^{*}$ satisfies weaker regularity conditions ($r=\tfrac{3}{4}$). In this case, T-kernel SGD also achieves the theoretically predicted convergence rate, while simultaneously attaining computational efficiency far superior to that of kernel SGD.
4.2 Robust Regression on 3-Dimensional Spherical Data
We further employ the three robust losses used in the previous subsection—Cauchy, Huber, and Welsch losses—to validate the main theoretical analysis on the three-dimensional sphere $\mathbb{S}^{2}$. Here, we consider the explanatory variable $X$ uniformly distributed on the sphere, with the response variable given by $Y=f^{*}(X)+\varepsilon$, where the noise term follows a Gaussian distribution $\varepsilon\sim\mathcal{N}(0,0.2^{2})$. The optimal fitting $f^{*}$ is defined as
$f^{*}=\frac{1}{5}\sum_{k=0}^{10}\left(\dim\Pi^{3}_{k}\right)^{-0.501-2sr}\sum_{j=1}^{2k+1}Y_{k,j},$
where $s=1$, $r=1$ and $\dim\Pi^{3}_{k}=\frac{(k+2)(k+1)}{2}+\frac{k(k+1)}{2}.$ In T-kernel SGD, we set the radius of the set $\mathcal{W}$ to $Q=1$, the step-size ratio to $\frac{\gamma_{n}}{\gamma_{0}}=n^{-\frac{2r}{2r+1}}$, the truncation parameter to $\theta=\frac{1}{2s(2r+1)}$, and adopt both the last iterate and the $\alpha$-suffix average with $\alpha=\tfrac{1}{2}$ as the output, consistent with the parameter setting in Theorem 1. In kernel SGD, we consider the Gaussian kernel and the Matérn-$\tfrac{5}{2}$ kernel from the previous subsection, as well as the following Matérn-$\tfrac{3}{2}$ kernel,
$K^{3/2}_{\mathrm{Matern}}(r)=\left(1+\sqrt{3}\,r\right)\exp\!\left(-\sqrt{3}\,r\right).$
We further set the step size in kernel SGD as $\gamma_{n}=\gamma_{0}n^{-\frac{2r}{2r+1}}$. The experimental results in Figure 3 demonstrate that T-kernel SGD achieves the theoretical optimality predicted in Theorem 1, while offering substantially higher computational efficiency compared to classical kernel SGD.
Figure 3: The left figure illustrates the convergence of the error with respect to the sample size under three different losses, while the right figure shows the convergence of the error with respect to runtime. The black line indicates the minimax rate, with the slope $\frac{2}{3}$.
4.3 Binary Classification of High-Dimensional Non-Spherical Data
In this subsection, we demonstrate the effective application of T-kernel SGD to real-world non-spherical datasets. Specifically, we employ the logistic loss to address the binary classification problem of distinguishing between even and odd digits in the MNIST dataset. The 784-dimensional MNIST dataset is widely used as a benchmark in machine learning to evaluate the performance of various algorithms. In the experiment, the output space $\mathcal{Y}=\{-1,1\}$ corresponds to the odd and even digits in the MNIST dataset, respectively. Additionally, we compare T-kernel SGD with the kernel SGD algorithm that utilizes a Gaussian kernel.
In T-kernel SGD, we define the inverse spherical-polar projection [31] as follows, which transforms non-spherical data into spherical data:
$F:\mathbb{R}^{d}_{+}$ $\to\mathbb{S}^{d},$
$x$ $\to\omega(x)=\frac{1}{4+x_{1}^{2}+\cdots+x_{d}^{2}}\left(4x_{1},\ldots,4x_{d},(4-x_{1}^{2}-\cdots-x_{d}^{2})\right).$
We select $K^{T}_{L_{n}}(x,x^{\prime})=\sum_{k=0}^{L_{n}}\left(\dim\Pi^{d}_{k}\right)^{-2s}K_{k}(x,x^{\prime})$ as the truncated kernel function in the recursive process, with the step size $\gamma_{n}=0.6n^{-0.05}$, and set the hyperparameters $\theta=0.68$ and $s=0.505$. For such real-world classification problems, the RKHS norm of the minimizer $f^{*}$ is unknown. Therefore, $Q$ is typically chosen sufficiently large; in this subsection, we set $Q=1000$. For convenience, we use the Polyak averaging and the last iterate as the output estimators. In the comparison experiment with kernel SGD, we use the standard Gaussian kernel $K(x,x^{\prime})=\exp\left(-\frac{\|x-x^{\prime}\|^{2}}{2\sigma^{2}}\right).$ Due to the high dimensionality of the data, a smoother Gaussian kernel with $\sigma=20$ is employed. Additionally, we apply Polyak averaging from [19] to enhance the robustness of the algorithm, and select a constant step size $\gamma_{n}=0.1$.
We augment the original MNIST dataset by adding Gaussian white noise. As shown in the sample-to-accuracy plot on the left, compared to kernel SGD, T-kernel SGD—despite applying gradient truncation—achieves better classification accuracy on the test dataset, demonstrating superior generalization performance. Meanwhile, the time-to-accuracy figure on the right further demonstrates that T-kernel SGD significantly improves computational efficiency, achieving a much higher accuracy than the classic kernel SGD within the same runtime. The numerical experiments above demonstrate that T-kernel SGD performs well on spherical, non-spherical, and datasets of varying dimensions.
Figure 4: The two plots above show the sample-accuracy and time-accuracy, respectively.
Appendix A Appendix
A.1 Preliminaries
In this section, we present the explicit expressions of spherical harmonics required for the algorithmic implementation, along with several auxiliary lemmas and their proofs used in the main text.
A.1.1 Orthonormal Basis of the Spherical Harmonic Space
[…]
$Y_{\alpha,0}$ $=h_{\alpha,0}\cdot g(x)\prod_{j=1}^{d-2}\left(x_{1}^{2}+\dots+x_{d-j+1}^{2}\right)^{\alpha_{j}/2}C_{\alpha_{j}}^{\lambda_{j}}\left(\frac{x_{d-j+1}}{\sqrt{x_{1}^{2}+\dots+x_{d-j+1}^{2}}}\right),\ \text{where }\alpha_{d-1}\geq 0,$
$Y_{\alpha,1}$ $=h_{\alpha,1}\cdot g(x)\prod_{j=1}^{d-2}\left(x_{1}^{2}+\dots+x_{d-j+1}^{2}\right)^{\alpha_{j}/2}C_{\alpha_{j}}^{\lambda_{j}}\left(\frac{x_{d-j+1}}{\sqrt{x_{1}^{2}+\dots+x_{d-j+1}^{2}}}\right),\ \text{where }\alpha_{d-1}\geq 1.$
Here, $h_{\alpha,i}$ with $i=0,1$ are normalization constants, and $C_{k}^{\lambda}(u)$ denotes the Gegenbauer polynomial, which satisfies $C_{0}^{\lambda}(u)=1$, $C_{1}^{\lambda}(u)=2\lambda u$, and the following three-term recurrence relation:
$C_{k+1}^{\lambda}(u)=\frac{2(k+\lambda)}{k+1}uC_{k}^{\lambda}(u)-\frac{k+2\lambda-1}{k+1}C_{k-1}^{\lambda}(u).$
For further properties of the Gegenbauer polynomials, we refer the reader to Appendix B.2 of [18]. For $Y_{\alpha,0}$, the function $g(x)$ corresponds to the real part of $(x_{2}+\sqrt{-1}\cdot x_{1})^{\alpha_{d-1}}$, whereas for $Y_{\alpha,1}$, $g(x)$ corresponds to the imaginary part of $(x_{2}+\sqrt{-1}\cdot x_{1})^{\alpha_{d-1}}$. The normalization constant $h_{\alpha,i}$ satisfies:
$h_{\alpha,i}^{-2}=\left\{\begin{aligned} &\frac{\pi}{\Omega_{d-1}}\prod_{j=1}^{d-2}\frac{\pi 2^{1-2\lambda_{j}}\Gamma(\alpha_{j}+2\lambda_{j})}{\alpha_{j}!(\lambda_{j}+\alpha_{j})\left(\Gamma(\lambda_{j})\right)^{2}},\ \ \text{if}\ \ \alpha_{d-1}>0,\\ &\frac{2\pi}{\Omega_{d-1}}\prod_{j=1}^{d-2}\frac{\pi 2^{1-2\lambda_{j}}\Gamma(\alpha_{j}+2\lambda_{j})}{\alpha_{j}!(\lambda_{j}+\alpha_{j})\left(\Gamma(\lambda_{j})\right)^{2}},\ \ \text{if}\ \ \alpha_{d-1}=0.\end{aligned}\right$.
Here, $\Gamma(u)$ represents the Gamma function. For a more detailed discussion of the orthonormal basis, we refer the reader to [18].
Next, we discuss the computational complexity of the basis functions. Since $0\leq\alpha_{j}\leq k$ and the quantities $\{x_{1}^{2},x_{1}^{2}+x_{2}^{2},\dots,x_{1}^{2}+\dots+x_{d}^{2}\}$ can be computed recursively, evaluating $\left(x_{1}^{2}+\dots+x_{d-j+1}^{2}\right)^{\alpha_{j}/2}$ requires at most $\mathcal{O}(k)$ computational time. Moreover, using the three-term recurrence relation of the Gegenbauer polynomials, computing $C_{\alpha_{j}}^{\lambda_{j}}\left(u\right)$ also requires at most $\mathcal{O}(k)$ computational time. Therefore, the computation of a basis function $Y_{\alpha,i}\in\mathcal{H}_{k}^{d}$ requires at most $\mathcal{O}(dk)$ computational time. Although the computational complexity of each basis function appears to increase with the data dimension, in high-dimensional settings (e.g., $d>100$) we typically only use second- or third-order polynomials. In such cases, the expressions of the orthonormal basis can be considerably simplified. The orthonormal bases for $\mathcal{H}_{0}^{d}$, $\mathcal{H}_{1}^{d}$, and $\mathcal{H}_{2}^{d}$ are given below
$\mathcal{H}_{0}^{d}=$ $\text{span}\{1\},$
$\mathcal{H}_{1}^{d}=$ $\text{span}\{\sqrt{d}x_{i}\}_{1\leq i\leq d},$
$\mathcal{H}_{2}^{d}=$ $\text{span}\Bigg(\left\{\sqrt{d(d+2)}x_{i}x_{j}\right\}_{1\leq i<j\leq d}\cup\left\{\frac{\sqrt{d(d+2)}}{2}(x_{1}^{2}-x_{2}^{2})\right\}$
$\quad\quad\cup\left\{h_{j}\cdot\left(x_{1}^{2}+\dots+x_{d-j+1}^{2}\right)C_{2}^{\lambda_{j}}\left(\frac{x_{d-j+1}}{\sqrt{x_{1}^{2}+\dots+x_{d-j+1}^{2}}}\right)\right\}_{1\leq j\leq d-2}\Bigg).$
The expression of the constant $h_{j}$ is given by:
$h_{j}=\pi^{1/2}\frac{(d-j)\left[(d-j)^{2}-1\right]}{d(d+2)}\frac{\Gamma(\frac{d-j+1}{2})}{\Gamma(\frac{d-j}{2})}\frac{2^{1-(d-j-1)}\Gamma(d-j-1)}{\left(\Gamma\left(\frac{d-j-1}{2}\right)\right)^{2}}.$
Since the Gamma function becomes computationally challenging in high dimensions, we consider simplifying the above expression using Poincaré-type expansions (see 5.11(i) in [42]) and the ratio of two Gamma functions (see [20]).
$2^{1-2\lambda}\frac{\Gamma(2\lambda)}{\left(\Gamma(\lambda)\right)^{2}}$ $=\frac{\lambda^{1/2}}{\pi^{1/2}}\frac{\Gamma^{*}(2\lambda)}{\left(\Gamma^{*}(\lambda)\right)^{2}},\quad\text{where}\ \ \Gamma^{*}(\lambda)=\sum_{k=0}^{\infty}\frac{g_{k}}{\lambda^{k}},$
$\frac{\Gamma\left(\lambda+\frac{1}{2}\right)}{\Gamma(\lambda)}$ $=\sum_{j=0}^{\infty}\frac{\left(2j-1-\frac{1}{2}\right)\left(2j-2-\frac{1}{2}\right)\dots\left(-\frac{1}{2}\right)}{(2j)!}B_{2j}^{(3/2)}\left(\frac{3}{4}\right)\left(\lambda-\frac{1}{4}\right)^{1/2-2j}.$
Here, $B_{2j}^{(3/2)}\left(\frac{3}{4}\right)$ represents the generalized Bernoulli polynomials, as detailed in [20].
A.1.2 Lemmas
Proposition A.1.
If $a_{k}>0$ and $\lim_{k\to\infty}a_{k}\cdot\left(\dim\Pi_{k}^{d}\right)^{2s}=l<\infty$ exists for some $s>\frac{1}{2}$, then the spherical radial basis function
$K(x,x^{\prime})=\sum_{k=0}^{\infty}a_{k}K_{k}(x,x^{\prime})$
defined in (2.3) converges uniformly and is uniformly bounded.
Proof.
[…]
$\dim{\mathcal{H}_{k}^{d}}=\dim{\mathcal{P}_{k}^{d}}-\dim{\mathcal{P}_{k-2}^{d}}=\binom{k+d-1}{d-1}-\binom{k+d-3}{d-1}$,
$\dim{\Pi_{k}^{d}}=\dim{\mathcal{P}_{k}^{d}}+\dim{\mathcal{P}_{k-1}^{d}}=\binom{k+d-1}{d-1}+\binom{k+d-2}{d-1}$.
Here $\dim\Pi_{k}^{d}$ satisfies the following relation
$\dim\Pi_{k}^{d}\geq\binom{k+d-1}{d-1}=\frac{(k+d-1)\dots(k+1)}{(d-1)!}\geq\frac{k^{d-1}}{(d-1)!}$
[…]
$|K(x,x^{\prime})|\leq\sum_{k=0}^{\infty}a_{k}|K_{k}(x,x^{\prime})|\leq M\left((d-1)!\right)^{2s}\sum_{k=0}^{\infty}k^{-2s(d-1)}\dim{\mathcal{H}_{k}^{d}}.$
If $d=2$, then $\dim\mathcal{H}_{k}^{d}=2$. In this case, when $s>\frac{1}{2}$, uniform convergence follows directly from the Weierstrass approximation theorem, since
$|K(x,x^{\prime})|\leq M\left((d-1)!\right)^{2s}\sum_{k=0}^{\infty}2k^{-2s}.$
If $d\geq 3$, then
$\dim{\mathcal{H}_{k}^{d}}=\binom{k+d-1}{d-1}-\binom{k+d-3}{d-1}$
$=$ $\frac{(k+d-1)(k+d-2)-k(k-1)}{(d-1)!}(k+d-3)\dots(k+1)$
$=$ $\frac{(d-1)(2k+d-2)(k+d-3)\dots(k+1)}{(d-1)!}\leq 2(k+1)^{d-2}.$
In this case, when $s>\frac{1}{2}$, uniform convergence follows directly from the Weierstrass approximation theorem, since
$|K(x,x^{\prime})|\leq M\left((d-1)!\right)^{2s}\sum_{k=0}^{\infty}2(k+1)^{d-2}k^{-2s(d-1)}\leq 2^{d-1}M\left((d-1)!\right)^{2s}\sum_{k=0}^{\infty}k^{-2s}.$
The proposition then follows. ∎
Before proving results related to the Fréchet derivative of the population risk, we first introduce a necessary preliminary.
Proposition A.2.
If Assumption 1 and Assumption 2 hold, then the losses $\ell(u,v)$ satisfies the following uniform condition with respect to its second argument $v$: for all $(u,v)\in(-B,B)\times\mathcal{Y}$, and $\forall\epsilon>0$, there exists $\delta_{\epsilon}>0$ such that for all $(u^{\prime},v)\in[-B,B]\times\mathcal{Y}$ with $|u-u^{\prime}|<\delta_{\epsilon}$, we have
$\left|\frac{\ell(u+u^{\prime},v)-\ell(u,v)}{u^{\prime}}-\partial_{u}\ell(u,v)\right|\leq\left|\partial_{u}\ell(u+\eta u^{\prime},v)-\partial_{u}\ell(u,v)\right|<\epsilon.$ (A.1)
Proof.
For any $\epsilon>0$, choose $\delta_{\epsilon}=\frac{\epsilon}{L}>0$. Then, for any $(u_{1},v),(u_{2},v)\in[-B,B]\times\mathcal{Y}$ such that $|u_{1}-u_{2}|<\delta_{\epsilon}$, we have $|\partial_{u}\ell(u_{1},v)-\partial_{u}\ell(u_{2},v)|\leq L|u_{1}-u_{2}|<\epsilon$. For any fixed $v\in\mathcal{Y}$, by the Lagrange mean value theorem, if $|u^{\prime}|<\delta_{\epsilon}$ and $u,u+u^{\prime}\in[-B,B]$, then there exists $\eta\in(0,1)$ such that
$\ell(u+u^{\prime},v)-\ell(u,v)=\partial_{u}\ell(u+\eta u^{\prime},v)u^{\prime},$
then we have
$\left|\frac{\ell(u+u^{\prime},v)-\ell(u,v)}{u^{\prime}}-\partial_{u}\ell(u,v)\right|\leq\left|\partial_{u}\ell(u+\eta u^{\prime},v)-\partial_{u}\ell(u,v)\right|\leq L|\eta u^{\prime}|<\epsilon.$
Lemma A.1.
If Assumption 1 holds and we choose $f\in\mathcal{W}$, then the Fréchet derivative of the population risk $\mathcal{E}(f)$ can be expressed as follows,
$\nabla\mathcal{E}(f)\big|_{\mathcal{H}_{K}}=\mathbb{E}\left[\partial_{u}\ell(f(X),Y)K(X,\cdot)\right].$
Proof.
By Proposition A.1, for any $\epsilon>0$, choose $h\in\mathcal{H}_{K}$ such that $\|h\|_{\infty}\leq\|h\|_{K}\kappa<\delta_{\epsilon}$ and $\|h\|_{K}\leq B\frac{1}{\kappa}-\|f\|_{K}$. Then, for any $Y\in\mathcal{Y}$, we have
$\left|\ell(f(X)+h(X),Y)-\ell(f(X),Y)-\partial_{u}\ell(f(X),Y)h(X)\right|<\epsilon|h(X)|$ (A.2)
$<$ $\epsilon\|h\|_{K}\|K(X,\cdot)\|_{K}<\epsilon\|h\|_{K}\kappa.$
Taking expectations on both sides of (A.2) and applying Jensen’s inequality, we obtain
$\left|\mathcal{E}(f+h)-\mathcal{E}(f)-\mathbb{E}\left[\partial_{u}\ell(f(X),Y)h(X)\right]\right|$
$=$ $\left|\mathbb{E}\left[\ell(f(X)+h(X),Y)-\ell(f(X),Y)-\partial_{u}\ell(f(X),Y)h(X)\right]\right|$
$\leq$ $\mathbb{E}\left[\left|\ell(f(X)+h(X),Y)-\ell(f(X),Y)-\partial_{u}\ell(f(X),Y)h(X)\right|\right]$
$\leq$ $\epsilon\|h\|_{K}\kappa.$
Using the reproducing property, one can obtain
$\mathcal{E}(f+h)-\mathcal{E}(f)-\mathbb{E}\left[\left\langle\partial_{u}\ell(f(X),Y)K(X,\cdot),h\right\rangle_{K}\right]$
$=$ $\mathcal{E}(f+h)-\mathcal{E}(f)-\left\langle\mathbb{E}\left[\partial_{u}\ell(f(X),Y)K(X,\cdot)\right],h\right\rangle_{K}=o(\|h\|_{K})$
Finally, by using the definition of the Fréchet derivative [15], we complete the proof
$\nabla\mathcal{E}(f)\big|_{\mathcal{H}_{K}}=\mathbb{E}\left[\partial_{u}\ell(f(X),Y)K(X,\cdot)\right].$
Lemma A.2.
If Assumption 1 holds and we choose $f\in\mathcal{W}\cap\mathcal{H}_{L_{n}}$ with $L_{n}\in{\mathbb{N}}$, then the Fréchet derivative of the population risk $\mathcal{E}(f)$ in the RKHS $(\mathcal{H}_{L_{n}},\langle\cdot,\cdot\rangle_{K})$ is given by
$\nabla\mathcal{E}(f)\big|_{\mathcal{H}_{L_{n}}}=\mathbb{E}\left[\partial_{u}\ell(f(X),Y)K_{L_{n}}^{T}(X,\cdot)\right].$
We also have
$P_{\mathcal{H}_{L_{n}}}\left(\partial_{u}\ell(f(X_{n}),Y_{n})K(X_{n},\cdot)\right)=\partial_{u}\ell(f(X_{n}),Y_{n})K_{L_{n}}^{T}(X_{n},\cdot).$
Proof.
Similar to Lemma A.1, by Proposition A.1, for any $\epsilon>0$, choose $h\in\mathcal{H}_{L_{n}}$ such that $\|h\|_{\infty}\leq\|h\|_{K}\kappa<\delta_{\epsilon}$ and $\|h\|_{K}\leq B\frac{1}{\kappa}-\|f\|_{K}$. Then, for any $Y\in\mathcal{Y}$, we have
$\left|\ell(f(X)+h(X),Y)-\ell(f(X),Y)-\partial_{u}\ell(f(X),Y)h(X)\right|<\epsilon|h(X)|<\epsilon\|h\|_{K}\kappa.$
Similar, we have
$\left|\mathcal{E}(f+h)-\mathcal{E}(f)-\mathbb{E}\left[\partial_{u}\ell(f(X),Y)h(X)\right]\right|$
$\leq$ $\mathbb{E}\left[\left|\ell(f(X)+h(X),Y)-\ell(f(X),Y)-\partial_{u}\ell(f(X),Y)h(X)\right|\right]$
$\leq$ $\epsilon\|h\|_{K}\kappa.$
Using the reproducing property, one can obtain
$\mathcal{E}(f+h)-\mathcal{E}(f)-\left\langle\mathbb{E}\left[\partial_{u}\ell(f(X),Y)K_{L_{n}}^{T}(X,\cdot)\right],h\right\rangle_{K}=o(\|h\|_{K}).$
By using the definition of the Fréchet derivative [15], we have
$\nabla\mathcal{E}(f)\big|_{\mathcal{H}_{L_{n}}}=\mathbb{E}\left[\partial_{u}\ell(f(X),Y)K_{L_{n}}^{T}(X,\cdot)\right].$
By the definition of the kernel function $K(x,x^{\prime})$, we have
$K(X_{n},\cdot)=\sum_{k=0}^{\infty}a_{k}K_{k}(X_{n},\cdot)=K_{L_{n}}^{T}(X,\cdot)+\sum_{k=L_{n}+1}^{\infty}a_{k}K_{k}(X_{n},\cdot).$
Finally, since each $K_{k}(X_{n},\cdot)\in\mathcal{H}_{k}^{d}\subset\mathcal{H}_{L_{n}}^{\perp}$ for $k\geq L_{n}+1$, it follows that $\sum_{k=L_{n}+1}^{\infty}a_{k}K_{k}(X_{n},\cdot)\in\mathcal{H}_{L_{n}}^{\perp}$. Therefore, the conclusion holds by the uniqueness of the orthogonal decomposition. ∎
Lemma A.3.
If Assumption 1 and Assumption 2 hold and the optimal function $f^{*}$ is an interior point of $\mathcal{W}$, i.e., $\|f^{*}\|_{K}<Q$, then for any $f\in\mathcal{W}$, we have
$\mathcal{E}(f)-\mathcal{E}(f^{*})\leq\frac{L}{2}\left\|f\circ F-f^{*}\circ F\right\|_{\rho_{X}}^{2},$
where $L$ is the Lipschitz constant defined in Assumption 2.
Proof.
Fix any $v\in\mathcal{Y}$, and define a function $l(u)=\ell(u,v)$ on [B,B][-B,B]. Then $l(u)$ is $L$-smooth and satisfies
$|l^{\prime}(u_{1})-l^{\prime}(u_{2})|=\left|\partial_{u}\ell(u_{1},v)-\partial_{u}\ell(u_{2},v)\right|\leq L|u_{1}-u_{2}|$
for $u_{1},u_{2}\in[-B,B]$. Then $l(u)$ satisfies quadratic upper bound in Theorem 2.1.5 of [41], i.e.
$l(u_{2})$ $\leq l(u_{1})+l^{\prime}(u_{1})(u_{1}-u_{2})+\frac{L}{2}(u_{1}-u_{2})^{2},$ (A.3)
$\Rightarrow\ \ \ell(u_{2},v)$ $\leq\ell(u_{1},v)+\partial_{u}\ell(u_{1},v)(u_{1}-u_{2})+\frac{L}{2}(u_{1}-u_{2})^{2}.$
In addition, by substituting $f\in\mathcal{W}$ and $f^{*}$ into (A.3) and taking expectations on both sides, we obtain
$\mathbb{E}\left[\ell(f\circ F(X),Y)\right]-\mathbb{E}\left[\ell(f^{*}\circ F(X),Y)\right]$
$\leq$ $\mathbb{E}\left[\partial_{u}\ell(f^{*}\circ F(X),Y)(f\circ F(X)-f^{*}\circ F(X))\right]+\frac{L}{2}\mathbb{E}\left[(f\circ F(X)-f^{*}\circ F(X))^{2}\right]$
$=$ $\left\langle\mathbb{E}\left[\partial_{u}\ell(f^{*}\circ F(X),Y)K(F(X),\cdot)\right],f-f^{*}\right\rangle_{K}+\frac{L}{2}\|f\circ F-f^{*}\circ F\|_{\rho_{X}}^{2}$
$=$ $\left\langle\nabla\mathcal{E}(f^{*})|_{\mathcal{H}_{K}},f-f^{*}\right\rangle_{K}+\frac{L}{2}\|f\circ F-f^{*}\circ F\|_{\rho_{X}}^{2}$
$\overset{\text{(i)}}{=}$ $\frac{L}{2}\|f\circ F-f^{*}\circ F\|_{\rho_{X}}^{2}.$
Since $f^{*}$ is an interior point and by Theorem 7.1-5 in [15], we have $\nabla\mathcal{E}(f^{*})|_{\mathcal{H}_{K}}=0$, which justifies equality (i). Therefore, the proof is complete by
$\mathcal{E}(f)-\mathcal{E}(f^{*})\leq\frac{L}{2}\left\|f\circ F-f^{*}\circ F\right\|_{\rho_{X}}^{2}.$
Lemma A.4.
Let $\mathcal{W}$ be defined as in (2.7), and denote by $P_{\mathcal{W}}:\mathcal{H}_{K}\rightarrow\mathcal{W}$ the projection operator onto $\mathcal{W}$. Then, for any $f\in\mathcal{H}_{L_{n}}$, we have $P_{\mathcal{W}}(f)\in\mathcal{H}_{L_{n}}\cap\mathcal{W}$.
Proof.
Note that the orthogonal complement of $\mathcal{H}_{L_{n}}$ in $\mathcal{H}_{K}$ is $\mathcal{H}_{L_{n}}^{\perp}$. Hence, the projection $P_{\mathcal{W}}(f)$ admits an orthogonal decomposition of the form $P_{\mathcal{W}}(f)=f_{1}+f_{2}$ with $f_{1}\in\mathcal{H}_{L_{n}}$ and $f_{2}\in\mathcal{H}_{L_{n}}^{\perp}$. If $f\in\mathcal{H}_{L_{n}}$ and $f_{2}\neq 0$, then one has
$\min_{g\in\mathcal{W}}\|g-f\|_{K}^{2}$ $=\|P_{\mathcal{W}}(f)-f\|_{K}^{2}$ (A.4)
$=\|(f_{1}+f_{2})-f\|_{K}^{2}=\|f_{1}-f\|_{K}^{2}+\|f_{2}\|_{K}^{2}>\|f_{1}-f\|_{K}^{2}.$
Since $\|f_{1}\|_{K}\leq\|P_{\mathcal{W}}(f)\|_{K}\leq Q$, it follows that $f_{1}\in\mathcal{W}$. This implies that (A.4) contradicts the definition of the projection operator $P_{\mathcal{W}}$, and hence $f_{2}=0$, which further implies $P_{\mathcal{W}}(f)\in\mathcal{H}_{L_{n}}$.
Lemma A.5.
If Assumption 1 and Assumption 3 holds, we have $\mathcal{E}(f)$ is convex function on convex set $\mathcal{W}$. For $f,g\in\mathcal{W}$, we have inequality
$\mathcal{E}(g)-\mathcal{E}(f)-\left\langle\nabla\mathcal{E}(f)\big|_{\mathcal{H}_{K}},g-f\right\rangle_{K}\geq\frac{\mu}{2}\|g\circ F-f\circ F\|_{\rho_{X}}^{2}.$
Proof.
For any $f,g\in\mathcal{W}$, the local $\mu$-strong convexity of $\ell(u,v)$ implies that
$\ \ \ \mathbb{E}\left[\ell(g\circ F(X),Y)-\ell(f\circ F(X),Y)-\partial_{u}\ell(f\circ F(X),Y)(g\circ F(X)-f\circ F(X))\right]$
$\geq\frac{\mu}{2}\mathbb{E}\left[(g\circ F(X)-f\circ F(X))^{2}\right]$
$\Rightarrow$ $\mathcal{E}(g)-\mathcal{E}(f)-\left\langle\mathbb{E}\left[\partial_{u}\ell(f\circ F(X),Y)K(F(X),\cdot)\right],g-f\right\rangle_{K}\geq\frac{\mu}{2}\|g\circ F-f\circ F\|_{\rho_{X}}^{2}$
$\Rightarrow$ $\mathcal{E}(g)-\mathcal{E}(f)-\left\langle\nabla\mathcal{E}(f)\big|_{\mathcal{H}_{K}},g-f\right\rangle_{K}\geq\frac{\mu}{2}\|g\circ F-f\circ F\|_{\rho_{X}}^{2}\geq 0.$
Thus, $\mathcal{E}(f)$ is convex by Section 7.12-1 in [15], and the proof is complete. ∎
Lemma A.6.
If $f\in\mathcal{H}_{L_{n}}$ and is represented as $f=\sum_{k=0}^{L_{n}}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}f_{k,j}Y_{k,j}$, then we have
$P_{\mathcal{W}}(f)=\left\{\begin{aligned} &\frac{Q}{\|f\|_{K}}f=\frac{Q}{\left(\sum_{k=0}^{L_{n}}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}a_{k}f_{k,j}^{2}\right)^{\frac{1}{2}}}f,\ \text{if}\ \|f\|_{K}>Q\\ &f,\ \text{if}\ \|f\|_{K}\leq Q.\end{aligned}\right$. (A.5)
Proof.
By the definition of the projection operator $P_{\mathcal{W}}$, we have
$P_{\mathcal{W}}(f)=\arg\min_{g\in\mathcal{W}}\|f-g\|_{K}^{2}.$
Furthermore, by Lemma A.4, we know that for any $f\in\mathcal{H}_{L_{n}}$, the projection $P_{\mathcal{W}}(f)\in\mathcal{H}_{L_{n}}$. Hence, the problem reduces to
$\min_{g\in\mathcal{H}_{L_{n}}}$ $\frac{1}{2}\|f-g\|_{K}^{2}$ (A.6)
s.t. $\frac{1}{2}\|g\|_{K}^{2}\leq\frac{1}{2}Q^{2}.$
Using the generalized Fourier expansions of $f$ and $g$ with respect to the orthonormal basis $\{Y_{k,j}\}$, we can transform (A.6) into the finite-dimensional convex optimization problem given in (A.7). If we assume $g=\sum_{k=0}^{L_{n}}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}g_{k,j}Y_{k,j}$, then
$\min_{g\in\mathcal{H}_{L_{n}}}$ $\frac{1}{2}\|f-g\|_{K}^{2}=\frac{1}{2}\sum_{k=0}^{L_{n}}a_{k}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}\left(g_{k,j}-f_{k,j}\right)^{2}$ (A.7)
s.t. $\frac{1}{2}\|g\|_{K}^{2}=\frac{1}{2}\sum_{k=0}^{L_{n}}a_{k}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}\left(g_{k,j}\right)^{2}\leq\frac{1}{2}Q^{2}.$
For $\lambda>0$, the Lagrangian corresponding to (A.7) is given by
$L(g,\lambda)=$ $\frac{1}{2}\|f-g\|_{K}^{2}+\frac{\lambda}{2}\left(\|g\|_{K}^{2}-Q^{2}\right)$
$=$ $\frac{1}{2}\sum_{k=0}^{L_{n}}a_{k}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}\left(g_{k,j}-f_{k,j}\right)^{2}+\frac{\lambda}{2}\left(\sum_{k=0}^{L_{n}}a_{k}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}\left(g_{k,j}\right)^{2}-Q^{2}\right).$
The KKT condition can be obtained as follows
$\left\{\begin{aligned} &\frac{\partial L}{\partial g_{k,j}}=a_{k}\left(g_{k,j}-f_{k,j}\right)+\lambda a_{k}g_{k,j}=0,\\ &\lambda\left(\sum_{k=0}^{L_{n}}a_{k}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}\left(g_{k,j}\right)^{2}-Q^{2}\right)=0,\\ &\frac{1}{2}\sum_{k=0}^{L_{n}}a_{k}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}\left(g_{k,j}\right)^{2}\leq\frac{1}{2}Q^{2}.\end{aligned}\right$.
Eventually, we conclude that if $\|f\|_{K}\leq Q$, then $P_{\mathcal{W}}(f)=f$; otherwise, if $\|f\|_{K}>Q$,
$P_{\mathcal{W}}(f)$ $=\sum_{k=0}^{L_{n}}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}\frac{f_{k,j}}{1+\lambda}Y_{k,j},$ (A.8)
$\frac{1}{1+\lambda}$ $=\frac{Q}{\left(\sum_{k=0}^{L_{n}}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}a_{k}f_{k,j}^{2}\right)^{\frac{1}{2}}}.$
Since the function $\frac{1}{2}\|f-g\|_{K}^{2}$ is strongly convex, the KKT point in (A.8) corresponds to the unique optimal solution. This completes the proof. ∎
Lemma A.7.
If Assumption 6 holds, then there exists a constant $0<b_{\rho}<B_{\rho}$ such that
$b_{\rho}\Omega_{d-1}\|f\|_{\omega}^{2}\leq\|f\circ F\|_{\rho_{X}}^{2}\leq B_{\rho}\Omega_{d-1}\|f\|_{\omega}^{2},\ \ \forall f\in\mathcal{H}_{K}.$
Proof.
[…]
$\int_{\Omega}|f\circ F|^{2}d\rho_{X}=\int_{\Omega}|f\circ F|^{2}\frac{d\rho_{X}}{d\lambda}d\lambda\leq B_{\rho}^{\prime}\int_{\Omega}|f\circ F|^{2}d\lambda$
$=$ $B_{\rho}^{\prime}\sum_{j=1}^{m}\int_{U_{j}}h_{j}|f\circ F|^{2}d\lambda=B_{\rho}^{\prime}\sum_{j=1}^{m}\int_{\phi(U_{j})}h_{j}\circ\phi_{j}^{-1}|f\circ F\circ\phi_{j}^{-1}|^{2}d\lambda$
$\overset{\text{(i)}}{=}$ $B_{\rho}^{\prime}\sum_{j=1}^{m}\int_{\psi_{j}(F(U_{j}))}h_{j}\circ F^{-1}\circ\psi_{j}^{-1}\cdot|f\circ\psi_{j}^{-1}|^{2}\cdot\left|\det\nabla\phi_{j}\circ F^{-1}\circ\psi_{j}^{-1}\right|dx$
[…]
Since both $f\circ F$ and $f$ are continuous, we do not distinguish between the Riemann and Lebesgue integrals in the proof of this lemma. Equality (i) follows from Theorem 3.7.1 in [11]. Moreover, because $F$ is a diffeomorphism, we have $\left|\det\nabla\phi_{j}\circ F^{-1}\circ\psi_{j}^{-1}\right|>0$ everywhere. Since each $h_{j}$ has compact support, the Jacobian determinant is bounded above and away from zero on the support of $h_{j}$. Together with the fact that the partition of unity $\{h_{j}\}_{1\leq j\leq m}$ consists of finite elements, the upper bound in (ii) follows. Equality (iii) follows directly from the definition of the partition of unity. The lower bound inequality can be established in a similar way. ∎
A.2 Proof of Theorem 2 (Strong Convergence)
In contrast to the order of presentation in the main text, we begin by proving the strong convergence guarantee of the T-kernel SGD. We then present the proof of Theorem 2 directly.
First, we defined the $f_{L_{n}}$ is the projection of $f^{*}$ in $(\mathcal{H}_{L_{n}},\left\langle\cdot,\cdot\right\rangle_{K})$.
$\left\|\hat{f}_{n}-f_{L_{n}}\right\|_{K}^{2}$
$=$ $\left\|P_{\mathcal{W}}\left(\hat{f}_{n-1}-\gamma_{n}\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)K_{L_{n}}^{T}(F(X_{n}),\cdot)\right)-f_{L_{n}}\right\|_{K}^{2}$
$\overset{\text{(i)}}{\leq}$ $\left\|\hat{f}_{n-1}-\gamma_{n}\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)K_{L_{n}}^{T}(F(X_{n}),\cdot)-f_{L_{n}}\right\|_{K}^{2}$
$=$ $\left\|\hat{f}_{n-1}-f_{L_{n}}\right\|_{K}^{2}-2\gamma_{n}\left\langle\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)K_{L_{n}}^{T}(F(X_{n}),\cdot),\hat{f}_{n-1}-f_{L_{n}}\right\rangle_{K}$
$+\gamma_{n}^{2}\left|\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)\right|^{2}\|K_{L_{n}}^{T}(F(X_{n}),\cdot)\|_{K}^{2}.$
In (i), we use the result $\|P_{\mathcal{W}}(f)-P_{\mathcal{W}}(g)\|_{K}\leq\|f-g\|_{K}$ for $f,g\in\mathcal{H}_{K}$, as stated in Section 4.3-1 of [15], where $\mathcal{W}$ is a closed convex subset of $\mathcal{H}_{K}$, and $f_{L_{n}}\in\mathcal{W}$. Since $\partial_{u}\ell(u,v)$ is continuous on the compact set $B,B]\times\mathcal{Y}$, it is bounded on this set. That is, there exists $M>0$ such that $\left|\partial_{u}\ell(u,v)\right|\leq M$. Moreover, using $\|\hat{f}_{n-1}\|_{\infty}\leq\|\hat{f}_{n-1}\|_{K}\kappa<B$, and the bound
[…]
we obtain
$\left|\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)\right|^{2}\|K_{L_{n}}^{T}(F(X_{n}),\cdot)\|_{K}^{2}\leq M^{2}\kappa^{2}:=M_{1}^{2},$
where we define $M_{1}^{2}:=M^{2}\kappa^{2}$. Therefore, one has
$\left\|\hat{f}_{n}-f_{L_{n}}\right\|_{K}^{2}-\left\|\hat{f}_{n-1}-f_{L_{n}}\right\|_{K}^{2}$ (A.9)
$\leq$ $-2\gamma_{n}\left\langle\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)K_{L_{n}}^{T}(F(X_{n}),\cdot),\hat{f}_{n-1}-f_{L_{n}}\right\rangle_{K}+\gamma_{n}^{2}M_{1}^{2}.$
Since $\mathcal{W}\cap\mathcal{H}_{L_{n}}$ is a bounded and closed subset of the finite-dimensional space $\mathcal{H}_{L_{n}}$, it is compact. Moreover, since $\mathcal{E}(f)$ is continuous on $\mathcal{W}$, it attains its minimum on the compact set $\mathcal{W}\cap\mathcal{H}_{L_{n}}$. That is, there exists $f_{L_{n}}^{*}=\arg\min_{f\in\mathcal{W}\cap\mathcal{H}_{L_{n}}}\mathcal{E}(f)$. Taking expectations on both sides of (A.9), we obtain
$\mathbb{E}\left[\left\|\hat{f}_{n}-f_{L_{n}}\right\|_{K}^{2}\right]$ (A.10)
$\leq$ $\mathbb{E}\left[\left\|\hat{f}_{n-1}-f_{L_{n}}\right\|_{K}^{2}\right]-2\gamma_{n}\mathbb{E}\left[\left\langle\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)K_{L_{n}}^{T}(F(X_{n}),\cdot),\hat{f}_{n-1}-f_{L_{n}}\right\rangle_{K}\right]+\gamma_{n}^{2}M_{1}^{2}.$
$=$ $\mathbb{E}\left[\left\|\hat{f}_{n-1}-f_{L_{n}}\right\|_{K}^{2}\right]-2\gamma_{n}\mathbb{E}\left[\left\langle\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)K_{L_{n}}^{T}(F(X_{n}),\cdot),\hat{f}_{n-1}-f_{L_{n}}^{*}\right\rangle_{K}\right]$
$-2\gamma_{n}\mathbb{E}\left[\left\langle\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)K_{L_{n}}^{T}(F(X_{n}),\cdot),f_{L_{n}}^{*}-f_{L_{n}}\right\rangle_{K}\right]+\gamma_{n}^{2}M_{1}^{2}.$
Next, we apply Lemma A.8 and Lemma A.9 to derive
$\mathbb{E}\left[\left\langle\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)K_{L_{n}}^{T}(F(X_{n}),\cdot),\hat{f}_{n-1}-f_{L_{n}}^{*}\right\rangle_{K}\right]$ (A.11)
$\geq$ $\frac{\mu}{2}\mathbb{E}\left[\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}^{2}\right].$
and
$-\mathbb{E}\left[\left\langle\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)K_{L_{n}}^{T}(F(X_{n}),\cdot),f_{L_{n}}^{*}-f_{L_{n}}\right\rangle_{K}\right]$ (A.12)
$\leq$ $L\cdot\mathbb{E}\left[\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\|_{\rho_{X}}\cdot\|f_{L_{n}}^{*}\circ F-f_{L_{n}}\circ F\|_{\rho_{X}}\right]+\frac{L}{2}\|f_{L_{n}}\circ F-f^{*}\circ F\|_{\rho_{X}}^{2}.$
We combine (A.11) and (A.12) to continue (A.10),
$\mathbb{E}\left[\left\|\hat{f}_{n}-f_{L_{n}}\right\|_{K}^{2}\right]$ (A.13)
$\leq$ $\mathbb{E}\left[\left\|\hat{f}_{n-1}-f_{L_{n}}\right\|_{K}^{2}\right]-\gamma_{n}\mu\mathbb{E}\left[\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}^{2}\right]+\gamma_{n}L\|f_{L_{n}}\circ F-f^{*}\circ F\|_{\rho_{X}}^{2}$
$+2\gamma_{n}L\cdot\mathbb{E}\left[\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\|_{\rho_{X}}\cdot\|f_{L_{n}}^{*}\circ F-f_{L_{n}}\circ F\|_{\rho_{X}}\right]+\gamma_{n}^{2}M_{1}^{2}$
$=$ $\mathbb{E}\left[\left\|\hat{f}_{n-1}-f_{L_{n}}\right\|_{K}^{2}\right]-\frac{\gamma_{n}\mu}{2}\mathbb{E}\left[\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}^{2}\right]+\gamma_{n}L\|f_{L_{n}}\circ F-f^{*}\circ F\|_{\rho_{X}}^{2}$
$+2\gamma_{n}L\cdot\mathbb{E}\bigg[\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}\cdot\bigg(\left\|f_{L_{n}}^{*}\circ F-f_{L_{n}}\circ F\right\|_{\rho_{X}}$
$-\frac{\mu}{4L}\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}\bigg)\bigg]+\gamma_{n}^{2}M_{1}^{2}.$
By Lemma A.10, we have
$\mathbb{E}\left[\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}\left(\left\|f_{L_{n}}^{*}\circ F-f_{L_{n}}\circ F\right\|_{\rho_{X}}-\frac{\mu}{4L}\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}\right)\right]$ (A.14)
$\leq\frac{4L}{\mu}\left\|f_{L_{n}}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}^{2}.$
Combining (A.14) with the preceding steps to continue from (A.13) yields
$\mathbb{E}\left[\left\|\hat{f}_{n}-f_{L_{n}}\right\|_{K}^{2}\right]$ (A.15)
$\leq$ $\mathbb{E}\left[\left\|\hat{f}_{n-1}-f_{L_{n}}\right\|_{K}^{2}\right]-\frac{\gamma_{n}\mu}{2}\mathbb{E}\left[\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}^{2}\right]+\gamma_{n}L\|f_{L_{n}}\circ F-f^{*}\circ F\|_{\rho_{X}}^{2}$
$+\frac{8L^{2}}{\mu}\gamma_{n}\left\|f_{L_{n}}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}^{2}+\gamma_{n}^{2}M_{1}^{2}.$
We use the following inequality in conjunction with (A.15)
$\mathbb{E}\left[\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}\circ F\right\|_{\rho_{X}}^{2}\right]\leq 2\mathbb{E}\left[\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}^{2}\right]+2\mathbb{E}\left[\left\|f_{L_{n}}^{*}\circ F-f_{L_{n}}\circ F\right\|_{\rho_{X}}^{2}\right]$
$\Rightarrow$ $-\mathbb{E}\left[\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}^{2}\right]\leq-\frac{1}{2}\mathbb{E}\left[\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}\circ F\right\|_{\rho_{X}}^{2}\right]+\left\|f_{L_{n}}^{*}\circ F-f_{L_{n}}\circ F\right\|_{\rho_{X}}^{2},$
to obtain
$\mathbb{E}\left[\left\|\hat{f}_{n}-f_{L_{n}}\right\|_{K}^{2}\right]$ (A.16)
$\leq$ $\mathbb{E}\left[\left\|\hat{f}_{n-1}-f_{L_{n}}\right\|_{K}^{2}\right]-\frac{\gamma_{n}\mu}{4}\mathbb{E}\left[\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}\circ F\right\|_{\rho_{X}}^{2}\right]+\frac{\gamma_{n}\mu}{2}\left\|f_{L_{n}}^{*}\circ F-f_{L_{n}}\circ F\right\|_{\rho_{X}}^{2}$
$+\gamma_{n}L\|f_{L_{n}}\circ F-f^{*}\circ F\|_{\rho_{X}}^{2}+\frac{8L^{2}}{\mu}\gamma_{n}\left\|f_{L_{n}}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}^{2}+\gamma_{n}^{2}M_{1}^{2}.$
We note that the orthogonal complement of $\mathcal{H}_{L_{n}}$ in $\mathcal{H}_{K}$ is $\mathcal{H}_{L_{n}}^{\perp}$. Since $\hat{f}_{n}-f_{L_{n}}\in\mathcal{H}_{L_{n}}$ and $f_{L_{n+1}}-f_{L_{n}}=(f_{L_{n+1}}-f^{*})-(f_{L_{n}}-f^{*})\in\mathcal{H}_{L_{n}}^{\perp}$, it follows that $\hat{f}_{n}-f_{L_{n}}$ is orthogonal to $f_{L_{n+1}}-f_{L_{n}}$. Therefore, we obtain
$\mathbb{E}\left[\left\|\hat{f}_{n}-f_{L_{n}}\right\|_{K}^{2}\right]=\mathbb{E}\left[\left\|\hat{f}_{n}-f_{L_{n+1}}\right\|_{K}^{2}\right]+\left\|f_{L_{n+1}}-f_{L_{n}}\right\|_{K}^{2}.$
Substituting the above equation back into (A.16), one can obtain
$\mathbb{E}\left[\left\|\hat{f}_{n}-f_{L_{n+1}}\right\|_{K}^{2}\right]$ (A.17)
$\leq$ $\mathbb{E}\left[\left\|\hat{f}_{n-1}-f_{L_{n}}\right\|_{K}^{2}\right]-\frac{\gamma_{n}\mu}{4}\mathbb{E}\left[\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}\circ F\right\|_{\rho_{X}}^{2}\right]$
$+\gamma_{n}\left(\frac{\mu}{2}+\frac{8L^{2}}{\mu}\right)\left\|f_{L_{n}}^{*}\circ F-f_{L_{n}}\circ F\right\|_{\rho_{X}}^{2}+\gamma_{n}L\|f_{L_{n}}\circ F-f^{*}\circ F\|_{\rho_{X}}^{2}$
$+\gamma_{n}^{2}M_{1}^{2}+\left\|f_{L_{n+1}}-f_{L_{n}}\right\|_{K}^{2}.$
In Lemma A.11, we show that if $f\in\mathcal{H}_{L_{n}}$,
$\|f\circ F\|_{\rho_{X}}^{2}\geq\frac{A^{2}_{2}}{A_{1}}\frac{b_{\rho}\Omega_{d-1}}{(2d)^{2s}}\,n^{-2\theta s}\|f\|_{K}^{2}.$ (A.18)
In Lemma A.12, we establish the following inequality
$\|f_{L_{n}}^{*}\circ F-f_{L_{n}}\circ F\|_{\rho_{X}}^{2}$ $\leq\frac{L}{\mu}B_{\rho}\Omega_{d-1}A_{1}^{2r}\|g^{*}\|_{\omega}^{2}\left(n+1\right)^{-4\theta sr},$ (A.19)
$\|f_{L_{n}}\circ F-f^{*}\circ F\|_{\rho_{X}}^{2}$ $\leq B_{\rho}\Omega_{d-1}A_{1}^{2r}\|g^{*}\|_{\omega}^{2}\left(n+1\right)^{-4\theta sr}.$
In combination with (A.18) and (A.19), we continue (A.17) to obtain
$\mathbb{E}\left[\left\|\hat{f}_{n}-f_{L_{n+1}}\right\|_{K}^{2}\right]$ (A.20)
$\leq$ $\left(1-\frac{A^{2}_{2}}{A_{1}}\frac{b_{\rho}\mu\Omega_{d-1}}{4(2d)^{2s}}\gamma_{n}n^{-2\theta s}\right)\mathbb{E}\left[\left\|\hat{f}_{n-1}-f_{L_{n}}\right\|_{K}^{2}\right]$
$+\gamma_{n}\left(\frac{\mu}{2}+\frac{8L^{2}}{\mu}\right)\frac{L}{\mu}B_{\rho}\Omega_{d-1}A_{1}^{2r}\|g^{*}\|_{\omega}^{2}\left(n+1\right)^{-4\theta sr}$
$+\gamma_{n}LB_{\rho}\Omega_{d-1}A_{1}^{2r}\|g^{*}\|_{\omega}^{2}\left(n+1\right)^{-4\theta sr}+\gamma_{n}^{2}M_{1}^{2}+\left\|f_{L_{n+1}}-f_{L_{n}}\right\|_{K}^{2}.$
We choose $t=\frac{2r}{2r+1}$, set the step size as $\gamma_{n}=\gamma_{0}n^{-t}\log(n+1)$, and also set $\theta=\frac{1}{2s(2r+1)}$. Under this parameter setting, we obtain the following two identities: $t=4\theta sr$ and $t+2\theta s=1$, as well as the inequality $(n+1)^{-4\theta sr}\leq\frac{\gamma_{n}}{\gamma_{0}\log(2)}$. We set the initial step size as $\gamma_{0}=c\frac{A_{1}4(2d)^{2s}}{A^{2}_{2}b_{\rho}\mu\Omega_{d-1}}$, where the constant $c$ satisfies $\frac{1}{\log(2)}\leq c\leq\frac{2}{\log(3)}$. Substituting the above constants and inequalities into (A.20), we obtain
$\mathbb{E}\left[\left\|\hat{f}_{n}-f_{L_{n+1}}\right\|_{K}^{2}\right]$
$\leq$ $\left(1-c\frac{\log(n+1)}{n}\right)\mathbb{E}\left[\left\|\hat{f}_{n-1}-f_{L_{n}}\right\|_{K}^{2}\right]+\left\|f_{L_{n+1}}-f_{L_{n}}\right\|_{K}^{2}$
$+\gamma_{n}^{2}\left[\left(\left(\frac{\mu}{2}+\frac{8L^{2}}{\mu}\right)\frac{L}{\mu}+L\right)B_{\rho}\Omega_{d-1}A_{1}^{2r}\|g^{*}\|_{\omega}^{2}\frac{1}{\gamma_{0}\log(2)}+M_{1}^{2}\right]$
$\overset{\text{(i)}}{\leq}$ $\left(1-c\frac{\log(n+1)}{n}\right)\mathbb{E}\left[\left\|\hat{f}_{n-1}-f_{L_{n}}\right\|_{K}^{2}\right]+\left\|f_{L_{n+1}}-f_{L_{n}}\right\|_{K}^{2}$
$+n^{-2t}\left(\log(n+1)\right)^{2}P^{2},$
In (i), we define the quantity $P^{2}=\gamma_{0}^{2}\left[\left(\left(\frac{\mu}{2}+\frac{8L^{2}}{\mu}\right)\frac{L}{\mu}+L\right)B_{\rho}\Omega_{d-1}A_{1}^{2r}\|g^{*}\|_{\omega}^{2}\frac{1}{\gamma_{0}\log(2)}+M_{1}^{2}\right]$.
Consider the function $h(u)=\frac{\log(u+1)}{u}$, which is monotonically decreasing for $u\geq 2$. In particular, we have $\left(1-c\frac{\log(n+1)}{n}\right)\geq 0$ for $n\geq 2$. Based on the recursive relation for $\hat{f}_{n}$, we have
$\mathbb{E}\left[\left\|\hat{f}_{n}-f_{L_{n+1}}\right\|_{K}^{2}\right]$ (A.21)
$\leq$ $\left(c\log(2)-1\right)\prod_{l=2}^{n}\left(1-c\frac{\log(l+1)}{l}\right)\left\|\hat{f}_{0}-f_{L_{1}}\right\|_{K}^{2}$
$+\sum_{k=1}^{n}\prod_{l=k+1}^{n}\left(1-c\frac{\log(l+1)}{l}\right)\left\|f_{L_{k}}-f_{L_{k+1}}\right\|_{K}^{2}$
$+\sum_{k=1}^{n}\prod_{l=k+1}^{n}\left(1-c\frac{\log(l+1)}{l}\right)k^{-2t}\left(\log(k+1)\right)^{2}P^{2}$
Here, we apply Lemma A.13 and Lemma A.14 to further derive from (A.21), from which we obtain
$\mathbb{E}\left[\left\|\hat{f}_{n}-f_{L_{n+1}}\right\|_{K}^{2}\right]$
$\leq$ $\left(2Q^{2}+2A_{1}^{2r-1}\|g^{*}\|_{\omega}^{2}\right)(n+1)^{-\frac{2r-1}{2r+1}}+(4r+2)P^{2}(\log(n+1))^{2}(n+1)^{-\frac{2r-1}{2r+1}}.$
Using the third inequality in Lemma A.12, we complete the proof of Theorem 2,
$\mathbb{E}\left[\left\|\hat{f}_{n}-f^{*}\right\|_{K}^{2}\right]=\mathbb{E}\left[\left\|\hat{f}_{n}-f_{L_{n+1}}\right\|_{K}^{2}\right]+\left\|f_{L_{n+1}}-f^{*}\right\|_{K}^{2}$
$\leq$ $\left(2Q^{2}+3A_{1}^{2r-1}\|g^{*}\|_{\omega}^{2}\right)(n+1)^{-\frac{2r-1}{2r+1}}+(4r+2)P^{2}(\log(n+1))^{2}(n+1)^{-\frac{2r-1}{2r+1}}.$
A.2.1 Technical Results
Lemma A.8.
If the assumptions in Theorem 1 hold and the quantity
$\mathbb{E}\left[\left\langle\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)K_{L_{n}}^{T}(F(X_{n}),\cdot),\hat{f}_{n-1}-f_{L_{n}}^{*}\right\rangle_{K}\right]$
is defined as in (A.10), then we have
$\mathbb{E}\left[\left\langle\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)K_{L_{n}}^{T}(F(X_{n}),\cdot),\hat{f}_{n-1}-f_{L_{n}}^{*}\right\rangle_{K}\right]\geq\frac{\mu}{2}\mathbb{E}\left[\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}^{2}\right].$
Proof.
By the local strong convexity of the loss function in Assumption 3, we have
$\ell(f_{L_{n}}^{*}\circ F(X_{n}),Y_{n})\geq$ $\ell(\hat{f}_{n-1}\circ F(X_{n}),Y_{n})+\frac{\mu}{2}(f_{L_{n}}^{*}\circ F(X_{n})-\hat{f}_{n-1}\circ F(X_{n}))^{2}$ (A.22)
$+\partial_{u}\ell(\hat{f}_{n-1}\circ F(X_{n}),Y_{n})(f_{L_{n}}^{*}\circ F(X_{n})-\hat{f}_{n-1}\circ F(X_{n})).$
Taking expectation on both sides of (A.22), one can obtain
$\mathcal{E}(f_{L_{n}}^{*})\overset{\text{(i)}}{\geq}\mathcal{E}(\hat{f}_{n-1})+\mathbb{E}\left[\left\langle\partial_{u}\ell(\hat{f}_{n-1}\circ F(X_{n}),Y_{n})K_{L_{n}}^{T}(F(X_{n}),\cdot),f_{L_{n}}^{*}-\hat{f}_{n-1}\right\rangle_{K}\right]$
$\quad\quad\quad\quad+\frac{\mu}{2}\mathbb{E}\left[\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}^{2}\right]$
$\Rightarrow$ $\mathbb{E}\left[\left\langle\partial_{u}\ell(\hat{f}_{n-1}\circ F(X_{n}),Y_{n})K_{L_{n}}^{T}(F(X_{n}),\cdot),\hat{f}_{n-1}-f_{L_{n}}^{*}\right\rangle_{K}\right]$
$\geq$ $\left(\mathcal{E}(\hat{f}_{n-1})-\mathcal{E}(f_{L_{n}}^{*})\right)+\frac{\mu}{2}\mathbb{E}\left[\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}^{2}\right]$
$\geq$ $\frac{\mu}{2}\mathbb{E}\left[\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}^{2}\right],$
where (i) follows from the fact that $\hat{f}_{n-1},f_{L_{n}}^{*}\in\mathcal{H}_{L_{n}}$ and $\left(\mathcal{H}_{L_{n}},\langle\cdot,\cdot\rangle_{K}\right)$ is a RKHS associated with the kernel $K_{L_{n}}^{T}(x,x^{\prime})$. This completes the proof. ∎
Lemma A.9.
If assumptions in Theorem 1 holds and
$\mathbb{E}\left[\left\langle\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)K_{L_{n}}^{T}(F(X_{n}),\cdot),f_{L_{n}}^{*}-f_{L_{n}}\right\rangle_{K}\right]$
is defined as in (A.10), we obtain
$-\mathbb{E}\left[\left\langle\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)K_{L_{n}}^{T}(F(X_{n}),\cdot),f_{L_{n}}^{*}-f_{L_{n}}\right\rangle_{K}\right]$
$\leq$ $L\cdot\mathbb{E}\left[\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\|_{\rho_{X}}\cdot\|f_{L_{n}}^{*}\circ F-f_{L_{n}}\circ F\|_{\rho_{X}}\right]+\frac{L}{2}\|f_{L_{n}}\circ F-f^{*}\circ F\|_{\rho_{X}}^{2}.$
Proof.
We begin by decomposing the following expression
$-\mathbb{E}\left[\left\langle\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)K_{L_{n}}^{T}(F(X_{n}),\cdot),f_{L_{n}}^{*}-f_{L_{n}}\right\rangle_{K}\right]$ (A.23)
$=$ $-\mathbb{E}\left[\left\langle\left(\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)-\partial_{u}\ell\left(f_{L_{n}}^{*}\circ F(X_{n}),Y_{n}\right)\right)K_{L_{n}}^{T}(F(X_{n}),\cdot),f_{L_{n}}^{*}-f_{L_{n}}\right\rangle_{K}\right]$
$-\mathbb{E}\left[\left\langle\partial_{u}\ell\left(f_{L_{n}}^{*}\circ F(X_{n}),Y_{n}\right)K_{L_{n}}^{T}(F(X_{n}),\cdot),f_{L_{n}}^{*}-f_{L_{n}}\right\rangle_{K}\right].$
Let $\mathcal{D}_{n-1}$ be the $\sigma$-field defined by $\mathcal{D}_{n-1}=\sigma\left((X_{1},Y_{1}),\dots,(X_{n-1},Y_{n-1})\right)$. Considering the first term in (A.23), one has
$-\mathbb{E}\left[\left\langle\left(\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)-\partial_{u}\ell\left(f_{L_{n}}^{*}\circ F(X_{n}),Y_{n}\right)\right)K_{L_{n}}^{T}(F(X_{n}),\cdot),f_{L_{n}}^{*}-f_{L_{n}}\right\rangle_{K}\right]$ (A.24)
$\overset{\text{(i)}}{\leq}$ $\mathbb{E}\left[\left|\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)-\partial_{u}\ell\left(f_{L_{n}}^{*}\circ F(X_{n}),Y_{n}\right)\right|\cdot\left|f_{L_{n}}^{*}\circ F(X_{n})-f_{L_{n}}\circ F(X_{n})\right|\right]$
$\overset{\text{(ii)}}{\leq}$ $L\cdot\mathbb{E}\left[\left|\hat{f}_{n-1}\circ F(X_{n})-f_{L_{n}}^{*}\circ F(X_{n})\right|\cdot\left|f_{L_{n}}^{*}\circ F(X_{n})-f_{L_{n}}\circ F(X_{n})\right|\right]$
$=$ $L\cdot\mathbb{E}\left[\mathbb{E}\left[\left|\hat{f}_{n-1}\circ F(X_{n})-f_{L_{n}}^{*}\circ F(X_{n})\right|\cdot\left|f_{L_{n}}^{*}\circ F(X_{n})-f_{L_{n}}\circ F(X_{n})\right|\ \big|\mathcal{D}_{n-1}\right]\right]$
$\overset{\text{(iii)}}{\leq}$ $L\cdot\mathbb{E}\left[\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}\cdot\left\|f_{L_{n}}^{*}\circ F-f_{L_{n}}\circ F\right\|_{\rho_{X}}\right],$
Here, (i) follows from the fact that $f_{L_{n}},f_{L_{n}}^{*}\in\mathcal{H}_{L_{n}}$ and that $\left(\mathcal{H}_{L_{n}},\langle\cdot,\cdot\rangle_{K}\right)$ is a RKHS associated with the kernel $K_{L_{n}}^{T}(x,x^{\prime})$. In (ii), we apply the local $L$-smoothness assumption stated in Assumption 2. In (iii), we use the Cauchy–Schwarz inequality.
Since $\mathcal{E}(f)$ is convex on $\mathcal{W}$ by Lemma A.5, and following Section 7.12-1 in [15], we analyze the second term in (A.23).
$-\mathbb{E}\left[\left\langle\partial_{u}\ell\left(f_{L_{n}}^{*}\circ F(X_{n}),Y_{n}\right)K_{L_{n}}^{T}(F(X_{n}),\cdot),f_{L_{n}}^{*}-f_{L_{n}}\right\rangle_{K}\right]$ (A.25)
$=$ $\left\langle\nabla\mathcal{E}(f_{L_{n}}^{*})\big|_{\mathcal{H}_{L_{n}}},f_{L_{n}}-f_{L_{n}}^{*}\right\rangle_{K}$
$\leq$ $\mathcal{E}(f_{L_{n}})-\mathcal{E}(f_{L_{n}}^{*})\leq\mathcal{E}(f_{L_{n}})-\mathcal{E}(f^{*})\overset{\text{(i)}}{\leq}\frac{L}{2}\|f_{L_{n}}\circ F-f^{*}\circ F\|_{\rho_{X}}^{2},$
where (i) is due to Lemma A.3. Finally, combining (A.24) and (A.25), we obtain the conclusion of the lemma
$-\mathbb{E}\left[\left\langle\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)K_{L_{n}}^{T}(F(X_{n}),\cdot),f_{L_{n}}^{*}\circ F-f_{L_{n}}\circ F\right\rangle_{K}\right]$
$\leq$ $L\cdot\mathbb{E}\left[\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\|_{\rho_{X}}\cdot\|f_{L_{n}}^{*}\circ F-f_{L_{n}}\circ F\|_{\rho_{X}}\right]+\frac{L}{2}\|f_{L_{n}}\circ F-f^{*}\circ F\|_{\rho_{X}}^{2}.$
Lemma A.10.
If the quantity in the first line of the following expression is defined as in equation (A.13), then we obtain
$\mathbb{E}\left[\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}\left(\left\|f_{L_{n}}^{*}\circ F-f_{L_{n}}\circ F\right\|_{\rho_{X}}-\frac{\mu}{4L}\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}\right)\right]$
$\leq\frac{4L}{\mu}\left\|f_{L_{n}}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}^{2}.$
Proof.
We define the following measurable set
$G=\left\{\left\|f_{L_{n}}^{*}\circ F-f_{L_{n}}\circ F\right\|_{\rho_{X}}-\frac{\mu}{4L}\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}\geq 0\right\},$
Meanwhile, the complement of $G$ is
$G^{c}=\left\{\left\|f_{L_{n}}^{*}\circ F-f_{L_{n}}\circ F\right\|_{\rho_{X}}-\frac{\mu}{4L}\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}<0\right\}.$
We then define the corresponding indicator functions $\mathcal{X}_{G}$ and $\mathcal{X}_{G^{c}}$, and decompose the original expression accordingly using these indicators, which yields
$\mathbb{E}\left[\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}\left(\left\|f_{L_{n}}^{*}\circ F-f_{L_{n}}\circ F\right\|_{\rho_{X}}-\frac{\mu}{4L}\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}\right)\right]$
$=$ $\mathbb{E}\left[\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}\left(\left\|f_{L_{n}}^{*}\circ F-f_{L_{n}}\circ F\right\|_{\rho_{X}}-\frac{\mu}{4L}\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}\right)\mathcal{X}_{G}\right]$
$+$ $\mathbb{E}\left[\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}\left(\left\|f_{L_{n}}^{*}\circ F-f_{L_{n}}\circ F\right\|_{\rho_{X}}-\frac{\mu}{4L}\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}\right)\mathcal{X}_{G^{c}}\right]$
$\leq$ $\mathbb{E}\left[\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}\left(\left\|f_{L_{n}}^{*}\circ F-f_{L_{n}}\circ F\right\|_{\rho_{X}}-\frac{\mu}{4L}\left\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}\right)\mathcal{X}_{G}\right]$
$\leq$ $\mathbb{E}\left[\|\hat{f}_{n-1}\circ F-f_{L_{n}}^{*}\circ F\|_{\rho_{X}}\|f_{L_{n}}^{*}\circ F-f_{L_{n}}\circ F\|_{\rho_{X}}\mathcal{X}_{G}\right]$
$\overset{\text{(i)}}{\leq}$ $\frac{4L}{\mu}\left\|f_{L_{n}}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}^{2}.$
Here, (i) follows from the definition of the set $G$. This completes the proof. ∎
Lemma A.11.
Suppose that Assumption 6 holds. For any $f\in\mathcal{H}_{L_{n}}$ with $L_{n}=\min\left\{k\,|\,\dim{\Pi_{k}^{d}}\geq n^{\theta}\right\}$, we have
$\|f\circ F\|_{\rho_{X}}^{2}\geq\frac{A^{2}_{2}}{A_{1}}\frac{b_{\rho}\Omega_{d-1}}{(2d)^{2s}}\,n^{-2\theta s}\|f\|_{K}^{2}$
Here, $A_{1}\geq A_{2}>0$ denote the upper and lower bounds of $a_{k}\cdot\left(\dim\Pi_{k}^{d}\right)^{2s}$ for all $k$, respectively, i.e.,
$A_{2}\left(\dim\Pi_{k}^{d}\right)^{-2s}\leq a_{k}\leq A_{1}\left(\dim\Pi_{k}^{d}\right)^{-2s}.$
Proof.
We choose $f\in\mathcal{H}_{L_{n}}$ and set $f=\sum_{k=0}^{L_{n}}\sum_{j=1}^{\dim\Pi_{k}^{d}}f_{k,j}Y_{k,j}$. Since $a_{k}>0$ and $\lim_{k\to\infty}a_{k}\cdot\left(\dim\Pi_{k}^{d}\right)^{2s}=l$, it follows that there exist constants $A_{1}\geq A_{2}>0$ such that $A_{2}\left(\dim\Pi_{k}^{d}\right)^{-2s}\leq a_{k}\leq A_{1}\left(\dim\Pi_{k}^{d}\right)^{-2s}$ and for any $p\geq k$, we have
$\frac{A^{2}_{2}}{A_{1}}\frac{\left(\dim\Pi_{p}^{d}\right)^{-2s}}{a_{k}}\leq\frac{A_{2}}{A_{1}}\frac{a_{p}}{a_{k}}\leq\frac{A_{2}\left(\dim\Pi_{p}^{d}\right)^{-2s}}{a_{k}}\leq\frac{A_{2}\left(\dim\Pi_{k}^{d}\right)^{-2s}}{a_{k}}\leq 1.$
Combining the above two inequality and Lemma A.7, we have
[…]
$=$ $b_{\rho}\Omega_{d-1}\sum_{k=0}^{L_{n}}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}f_{k,j}^{2}\geq b_{\rho}\Omega_{d-1}\frac{A^{2}_{2}}{A_{1}}\left(\dim\Pi_{L_{n}}^{d}\right)^{-2s}\sum_{k=0}^{L_{n}}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}\frac{f_{k,j}^{2}}{a_{k}}$
$\overset{\text{(i)}}{\geq}$ $\frac{A^{2}_{2}}{A_{1}}\frac{b_{\rho}\Omega_{d-1}}{(2d)^{2s}}\,n^{-2\theta s}\|f\|_{K}^{2}$
In (i), we use $\dim\Pi_{L_{n}-1}^{d}\leq n^{\theta}\leq\dim\Pi_{L_{n}}^{d}$ and $\dim\Pi_{L_{n}}^{d}\leq 2d\cdot\dim\Pi_{L_{n}-1}^{d}$ in Lemma 12 in [5], where we defined $\dim\Pi_{-1}^{d}=1$.
Lemma A.12.
If assumptions in Theorem 1 holds, for $L_{m}\geq L_{n}\in{\mathbb{N}}$, we have
$\|f_{L_{n}}^{*}\circ F-f_{L_{n}}\circ F\|_{\rho_{X}}^{2}$ $\leq\frac{L}{\mu}B_{\rho}\Omega_{d-1}A_{1}^{2r}\|g^{*}\|_{\omega}^{2}\left(n+1\right)^{-4\theta sr},$
$\|f_{L_{n}}\circ F-f^{*}\circ F\|_{\rho_{X}}^{2}$ $\leq B_{\rho}\Omega_{d-1}A_{1}^{2r}\|g^{*}\|_{\omega}^{2}\left(n+1\right)^{-4\theta sr}$
and we also have
$\left\|f_{L_{n}}-f^{*}\right\|_{K}^{2}$ $\leq A_{1}^{2r-1}\left(n+1\right)^{-2\theta s(2r-1)}\|g^{*}\|_{\omega}^{2},$
$\left\|f_{L_{n}}-f_{L_{m}}\right\|_{K}^{2}$ $\leq A_{1}^{2r-1}\left(n+1\right)^{-2\theta s(2r-1)}\|g^{*}\|_{\omega}^{2}.$
Proof.
First, we use Locally $\mu$-strong convex to obtain
$\mathcal{L}(f_{L_{n}}\circ F(X_{n}),Y_{n})\geq$ $\mathcal{L}(f_{L_{n}}^{*}\circ F(X_{n}),Y_{n})$ (A.26)
$+\partial_{u}\ell(f_{L_{n}}^{*}\circ F(X_{n}),Y_{n})(f_{L_{n}}\circ F(X_{n})-f_{L_{n}}^{*}\circ F(X_{n}))$
$+\frac{\mu}{2}(f_{L_{n}}^{*}\circ F(X_{n})-f_{L_{n}}\circ F(X_{n}))^{2},$
Taking expectation on both sides of (A.26) to yield
$\mathcal{E}(f_{L_{n}})-\mathcal{E}(f_{L_{n}}^{*})$ (A.27)
$\geq$ $\mathbb{E}\left[\left\langle\partial_{u}\ell(f_{L_{n}}^{*}\circ F(X_{n}),Y_{n})K_{L_{n}}^{T}(F(X_{n}),\cdot),f_{L_{n}}-f_{L_{n}}^{*}\right\rangle_{K}\right]+\frac{\mu}{2}\left\|f_{L_{n}}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}^{2}$
$\overset{\text{(i)}}{\geq}$ $\frac{\mu}{2}\left\|f_{L_{n}}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}^{2}.$
Here, (i) follows from the Euler inequality of the convex function $\mathcal{E}(f)$ at its minimizer $f_{L_{n}}^{*}$ over the convex set $\mathcal{W}\cap\mathcal{H}_{L_{n}}$ (see Lemma A.5 and Theorem 7.12-3 in [15]). Then by Lemma A.3, we using (A.27) to obtain
$\frac{\mu}{2}\left\|f_{L_{n}}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}^{2}\leq\mathcal{E}(f_{L_{n}})-\mathcal{E}(f_{L_{n}}^{*})\leq\mathcal{E}(f_{L_{n}})-\mathcal{E}(f^{*})\leq\frac{L}{2}\left\|f_{L_{n}}\circ F-f^{*}\circ F\right\|_{\rho_{X}}^{2}.$ (A.28)
Following a similar argument as in the proof of Lemma A.11, for $k\geq l$, we have
$1\leq A_{1}\frac{\left(\dim\Pi_{k}^{d}\right)^{-2s}}{a_{k}}\leq A_{1}\frac{\left(\dim\Pi_{l}^{d}\right)^{-2s}}{a_{k}}.$
Let us denote $f^{*}=\sum_{k=0}^{\infty}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}f_{k,j}^{*}Y_{k,j}$. By applying Lemma A.7, we obtain
$\left\|f_{L_{n}}\circ F-f^{*}\circ F\right\|_{\rho_{X}}^{2}\leq$ $B_{\rho}\Omega_{d-1}\|f_{L_{n}}-f^{*}\|_{\omega}^{2}$ (A.29)
[…]
$=$ $B_{\rho}\Omega_{d-1}\sum_{k=L_{n}+1}^{\infty}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}\left(f_{k,j}^{*}\right)^{2}$
$\leq$ $B_{\rho}\Omega_{d-1}A_{1}^{2r}\left(\dim\Pi_{L_{n}+1}^{d}\right)^{-4sr}\sum_{k=L_{n}+1}^{\infty}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}\frac{f_{k,j}^{2}}{a_{k}^{2r}}$
$\overset{\text{(i)}}{\leq}$ $B_{\rho}\Omega_{d-1}A_{1}^{2r}\|g^{*}\|_{\omega}^{2}\left(n+1\right)^{-4\theta sr}.$
In (i), we use $(n+1)^{\theta}\leq\dim\Pi_{L_{n+1}}^{d}\leq\dim\Pi_{L_{n}+1}^{d}$. Combining (A.28) and (A.29), one has
$\left\|f_{L_{n}}\circ F-f_{L_{n}}^{*}\circ F\right\|_{\rho_{X}}^{2}\leq\frac{L}{\mu}\left\|f_{L_{n}}\circ F-f^{*}\circ F\right\|_{\rho_{X}}^{2}\leq\frac{L}{\mu}B_{\rho}\Omega_{d-1}A_{1}^{2r}\|g^{*}\|_{\omega}^{2}\left(n+1\right)^{-4\theta sr}.$
Next we prove the last two inequalities,
$\left\|f_{L_{n}}-f^{*}\right\|_{K}^{2}$ $=\sum_{k=L_{n}+1}^{\infty}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}\frac{(f_{k,j}^{*})^{2}}{a_{k}}\leq\sum_{k=L_{n}+1}^{\infty}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}\frac{(f_{k,j}^{*})^{2}}{a_{k}}A_{1}^{2r-1}\frac{\left(\dim\Pi_{L_{n}+1}^{d}\right)^{-2s(2r-1)}}{a_{k}^{2r-1}}$
$\leq A_{1}^{2r-1}\left(\dim\Pi_{L_{n}+1}^{d}\right)^{-2s(2r-1)}\sum_{k=L_{n}+1}^{\infty}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}\frac{(f_{k,j}^{*})^{2}}{a_{k}^{2r}}$
$\leq A_{1}^{2r-1}\left(n+1\right)^{-2\theta s(2r-1)}\|g^{*}\|_{\omega}^{2},$
and
$\left\|f_{L_{n}}-f_{L_{m}}\right\|_{K}^{2}$ $\leq\left\|f_{L_{n}}-f_{L_{m}}\right\|_{K}^{2}+\left\|f_{L_{m}}-f^{*}\right\|_{K}^{2}$
$=\left\|f_{L_{n}}-f^{*}\right\|_{K}^{2}\leq A_{1}^{2r-1}\left(n+1\right)^{-2\theta s(2r-1)}\|g^{*}\|_{\omega}^{2}.$
The proof is now complete. ∎
Lemma A.13.
If $\frac{1}{\log(2)}\leq c\leq\frac{2}{\log(3)}$ and $t=\frac{2r}{2r+1}$, then we have
$\sum_{k=1}^{n}\prod_{l=k+1}^{n}\left(1-c\frac{\log(l+1)}{l}\right)k^{-2t}\left(\log(k+1)\right)^{2}\leq(4r+2)(\log(n+1))^{2}(n+1)^{-\frac{2r-1}{2r+1}}.$
Proof.
Since $\frac{1}{\log(2)}\leq c\leq\frac{2}{\log(3)}$, it follows that $0\leq\left(1-c\frac{\log(l+1)}{l}\right)\leq\left(1-\frac{1}{l}\right)=\frac{l-1}{l}$ for all $l\geq 2$. We can then obtain
$\sum_{k=1}^{n}\prod_{l=k+1}^{n}\left(1-c\frac{\log(l+1)}{l}\right)k^{-2t}\left(\log(k+1)\right)^{2}$
$\leq$ $\sum_{k=1}^{n}\prod_{l=k+1}^{n}\left(1-\frac{1}{l}\right)k^{-2t}\left(\log(k+1)\right)^{2}$
$\leq$ $\left(\log(n+1)\right)^{2}\sum_{k=1}^{n}\left(\prod_{l=k+1}^{n}\frac{l-1}{l}\right)k^{-2t}$
$=$ $\left(\log(n+1)\right)^{2}\frac{1}{n}\sum_{k=1}^{n}k^{-2t+1}\leq 4\left(\log(n+1)\right)^{2}\frac{1}{n+1}\sum_{k=1}^{n}(k+1)^{-2t+1}$
$\leq$ $4\left(\log(n+1)\right)^{2}\frac{1}{n+1}\int_{1}^{n+1}x^{1-2t}dx\leq\frac{2\left(\log(n+1)\right)^{2}}{(n+1)(1-t)}(n+1)^{2-2t}$
$=$ $(4r+2)\left(\log(n+1)\right)^{2}(n+1)^{-\frac{2r-1}{2r+1}}.$
This completes the proof. ∎
Lemma A.14.
If the assumptions in Theorem 1 hold, we have
$\left(c\log(2)-1\right)\prod_{l=2}^{n}\left(1-c\frac{\log(l+1)}{l}\right)\left\|\hat{f}_{0}-f_{L_{1}}\right\|_{K}^{2}$ (A.30)
$+\sum_{k=1}^{n}\prod_{l=k+1}^{n}\left(1-c\frac{\log(l+1)}{l}\right)\left\|f_{L_{k}}-f_{L_{k+1}}\right\|_{K}^{2}$
$\leq$ $\left(2Q^{2}+2A_{1}^{2r-1}\|g^{*}\|_{\omega}^{2}\right)(n+1)^{-\frac{2r-1}{2r+1}}.$
Proof.
First, we consider the second term in (A.30)
$\sum_{k=1}^{n}\prod_{l=k+1}^{n}\left(1-c\frac{\log(l+1)}{l}\right)\left\|f_{L_{k}}-f_{L_{k+1}}\right\|_{K}^{2}$ (A.31)
$\leq$ $\sum_{k=1}^{\frac{n}{2}-\frac{1}{2}}\prod_{l=k+1}^{n}\left(1-c\frac{\log(l+1)}{l}\right)\left\|f_{L_{k}}-f_{L_{k+1}}\right\|_{K}^{2}$
$+\sum_{k=\frac{n}{2}-\frac{1}{2}}^{n}\prod_{l=k+1}^{n}\left(1-c\frac{\log(l+1)}{l}\right)\left\|f_{L_{k}}-f_{L_{k+1}}\right\|_{K}^{2}$
$\overset{\text{(i)}}{\leq}$ $\sum_{k=1}^{\frac{n}{2}-\frac{1}{2}}\prod_{l=k+1}^{n}\left(1-c\frac{\log(l+1)}{l}\right)\left\|f_{L_{k}}-f_{L_{k+1}}\right\|_{K}^{2}+\sum_{k=\frac{n}{2}-\frac{1}{2}}^{n}\left\|f_{L_{k}}-f_{L_{k+1}}\right\|_{K}^{2}$
$\overset{\text{(ii)}}{=}$ $\sum_{k=1}^{\frac{n}{2}-\frac{1}{2}}\prod_{l=k+1}^{n}\left(1-c\frac{\log(l+1)}{l}\right)\left\|f_{L_{k}}-f_{L_{k+1}}\right\|_{K}^{2}+\left\|f_{L_{\frac{n}{2}-\frac{1}{2}}}-f_{L_{n+1}}\right\|_{K}^{2},$
$=$ $\prod_{l=2}^{n}\left(1-c\frac{\log(l+1)}{l}\right)\left\|f_{L_{1}}-f_{L_{2}}\right\|_{K}^{2}+\sum_{k=2}^{\frac{n}{2}-\frac{1}{2}}\prod_{l=k+1}^{n}\left(1-c\frac{\log(l+1)}{l}\right)\left\|f_{L_{k}}-f_{L_{k+1}}\right\|_{K}^{2}$
$+\left\|f_{L_{\frac{n}{2}-\frac{1}{2}}}-f_{L_{n+1}}\right\|_{K}^{2}.$
Here, (i) follows from the inequality $\frac{1}{\log(2)}\leq c\leq\frac{2}{\log(3)}$, which implies that $0\leq\left(1-c\frac{\log(l+1)}{l}\right)\leq\left(1-\frac{1}{l}\right)\leq 1$ for all $l\geq 2$. Consider the two terms $f_{L_{m+1}}-f_{L_{m}}$ and $f_{L_{k+1}}-f_{L_{k}}$ for indices $m>k$. The difference $f_{L_{k+1}}-f_{L_{k}}$ belongs to $\mathcal{H}_{L_{k}}$, while the difference $f_{L_{m+1}}-f_{L_{m}}=(f_{L_{m+1}}-f^{*})-(f_{L_{m}}-f^{*})$ lies in the orthogonal complement $\mathcal{H}_{L_{k}}^{\perp}$. Therefore, $f_{L_{m+1}}-f_{L_{m}}$ and $f_{L_{k+1}}-f_{L_{k}}$ are orthogonal, and condition (ii) is satisfied.
Since $\hat{f}_{0}=0$, we now bound the first terms in both (A.30) and (A.31),
$\left(c\log(2)-1\right)\prod_{l=2}^{n}\left(1-c\frac{\log(l+1)}{l}\right)\left\|\hat{f}_{0}-f_{L_{1}}\right\|_{K}^{2}+\prod_{l=2}^{n}\left(1-c\frac{\log(l+1)}{l}\right)\left\|f_{L_{1}}-f_{L_{2}}\right\|_{K}^{2}$ (A.32)
$\leq$ $\left(c\log(2)-1\right)\prod_{l=2}^{n}\left(1-\frac{1}{l}\right)\left\|\hat{f}_{0}-f_{L_{1}}\right\|_{K}^{2}+\prod_{l=2}^{n}\left(1-\frac{1}{l}\right)\left\|f_{L_{1}}-f_{L_{2}}\right\|_{K}^{2}$
$\leq$ $\frac{1}{n}\left\|f_{L_{1}}\right\|_{K}^{2}+\frac{1}{n}\left\|f_{L_{1}}-f_{L_{2}}\right\|_{K}^{2}=\frac{1}{n}\left\|f_{L_{2}}\right\|_{K}^{2}.$
For $2\leq k\leq\frac{n}{2}-\frac{1}{2}$, we have
$\prod_{l=k+1}^{n}\left(1-c\frac{\log(l+1)}{l}\right)\leq$ $\exp\left(\sum_{l=k+1}^{n}\log\left(1-c\frac{\log(l+1)}{l}\right)\right)$ (A.33)
$\leq$ $\exp\left(-c\sum_{l=k+1}^{n}\frac{\log(l+1)}{l}\right)\leq\exp\left(-c\sum_{l=k+1}^{n}\frac{\log(l)}{l}\right)$
$\overset{\text{(i)}}{\leq}$ $\exp\left(-c\int_{x=k+1}^{n+1}\frac{\log(x)}{x}dx\right)$
$=$ $\exp\left(-\frac{c}{2}\left[\left(\log(n+1)\right)^{2}-\left(\log\left(k+1\right)\right)^{2}\right]\right)$
$\leq$ $\exp\left(-\frac{c}{2}\left[\left(\log(n+1)\right)^{2}-\left(\log\left(\frac{n+1}{2}\right)\right)^{2}\right]\right)$
$\leq$ $\exp\left(-\frac{c}{2}\left[\left(\log(n+1)\right)^{2}-\left(\log\left(n+1\right)-\log(2)\right)^{2}\right]\right)$
$=$ $\exp\left(\frac{c}{2}\left(\log(2)\right)^{2}\right)\exp\left(-c\log(2)\log(n+1)\right)$
$\leq$ $2\exp\left(-c\log(2)\log(n+1)\right)=\frac{2}{(n+1)^{c\log(2)}}$
$\overset{\text{(ii)}}{\leq}$ $\frac{2}{n+1}.$
The function $\frac{\log(x)}{x}$ has derivative $\frac{1-\log(x)}{x^{2}}$, so it is decreasing for $x\geq e$. Thus, the inequality in (i) holds. In (ii), we use the inequality $\frac{1}{\log(2)}\leq c\leq\frac{2}{\log(3)}$. Next, we return to the second term in (A.31). By incorporating (A.33), we then obtain
$\sum_{k=2}^{\frac{n}{2}-\frac{1}{2}}\prod_{l=k+1}^{n}\left(1-c\frac{\log(l+1)}{l}\right)\left\|f_{L_{k}}-f_{L_{k+1}}\right\|_{K}^{2}$ (A.34)
$\leq$ $\frac{2}{n+1}\sum_{k=2}^{\frac{n}{2}-\frac{1}{2}}\left\|f_{L_{k}}-f_{L_{k+1}}\right\|_{K}^{2}=\frac{2}{n+1}\left\|f_{L_{2}}-f_{L_{\frac{n+1}{2}}}\right\|_{K}^{2}.$
Finally, substituting the estimates from (A.32) and (A.34) into (A.31) yields
$\left(c\log(2)-1\right)\prod_{l=2}^{n}\left(1-c\frac{\log(l+1)}{l}\right)\left\|\hat{f}_{0}-f_{L_{1}}\right\|_{K}^{2}$
$+\sum_{k=1}^{n}\prod_{l=k+1}^{n}\left(1-c\frac{\log(l+1)}{l}\right)\left\|f_{L_{k}}-f_{L_{k+1}}\right\|_{K}^{2}$
$\leq$ $\left(c\log(2)-1\right)\prod_{l=2}^{n}\left(1-c\frac{\log(l+1)}{l}\right)\left\|\hat{f}_{0}-f_{L_{1}}\right\|_{K}^{2}+\prod_{l=2}^{n}\left(1-c\frac{\log(l+1)}{l}\right)\left\|f_{L_{1}}-f_{L_{2}}\right\|_{K}^{2}$
$+\sum_{k=2}^{\frac{n}{2}-\frac{1}{2}}\prod_{l=k+1}^{n}\left(1-c\frac{\log(l+1)}{l}\right)\left\|f_{L_{k}}-f_{L_{k+1}}\right\|_{K}^{2}+\left\|f_{L_{\frac{n}{2}-\frac{1}{2}}}-f_{L_{n+1}}\right\|_{K}^{2}$
$\leq$ $\frac{1}{n}\left\|f_{L_{2}}\right\|_{K}^{2}+\frac{2}{n+1}\left\|f_{L_{2}}-f_{L_{\frac{n+1}{2}}}\right\|_{K}^{2}+\left\|f_{L_{\frac{n}{2}-\frac{1}{2}}}-f_{L_{n+1}}\right\|_{K}^{2}$
$\leq$ $\frac{2}{n+1}\left\|f_{L_{\frac{n+1}{2}}}\right\|_{K}^{2}+\left\|f_{L_{\frac{n}{2}-\frac{1}{2}}}-f_{L_{n+1}}\right\|_{K}^{2}$
$\leq$ $\frac{2}{n+1}\left\|f^{*}\right\|_{K}^{2}+\left\|f_{L_{\frac{n}{2}-\frac{1}{2}}}-f_{L_{n+1}}\right\|_{K}^{2}$
$\overset{\text{(i)}}{\leq}$ $\frac{2}{n+1}Q^{2}+A_{1}^{2r-1}\left(\frac{n+1}{2}\right)^{-2\theta s(2r-1)}\|g^{*}\|_{\omega}^{2}$
$\leq$ $\left(2Q^{2}+2A_{1}^{2r-1}\|g^{*}\|_{\omega}^{2}\right)(n+1)^{-\frac{2r-1}{2r+1}}.$
Here, (i) follows from the Assumption 5 that $f^{*}\in\mathcal{W}=\left\{f\in\mathcal{H}_{K}\,|\ \|f\|_{K}\leq Q\right\}$ and from the inequality $\left\|f_{L_{n}}-f_{L_{m}}\right\|_{K}^{2}\leq A_{1}^{2r-1}\left(n+1\right)^{-2\theta s(2r-1)}\|g^{*}\|_{\omega}^{2}$ for $L_{m}\geq L_{n}\in\mathbb{N}$, as stated in Lemma A.12. This completes the proof. ∎
A.3 Proof of Theorem 1
In this section, we use the result of Theorem 2 to prove the main result of the paper, Theorem 1. We begin by analyzing the convergence of the $\alpha$-suffix average $\bar{f}_{\alpha n}$.
A.3.1 Convergence Analysis of Suffix Averaging
Let the constant be $\widetilde{C}=\left[\frac{\left(2Q^{2}+3A_{1}^{2r-1}\|g^{*}\|_{\omega}^{2}\right)}{(\log(2))^{2}}+(4r+2)P^{2}\right]$. Then, the convergence result in Theorem 2 can be rewritten as follows
$\mathbb{E}\left[\left\|\hat{f}_{n}-f^{*}\right\|_{K}^{2}\right]$ $\leq\widetilde{C}\left(\log(n+1)\right)^{2}(n+1)^{-\frac{2r-1}{2r+1}}.$
Based on the recursive formula of $\hat{f}_{n}$ in (2.11), we obtain
$\mathbb{E}\left[\left\|\hat{f}_{n}-f^{*}\right\|_{K}^{2}\right]$ (A.35)
$=$ $\mathbb{E}\left[\left\|P_{\mathcal{W}}\left(\hat{f}_{n-1}-\gamma_{n}\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)K_{L_{n}}^{T}(F(X_{n}),\cdot)\right)-f^{*}\right\|_{K}^{2}\right]$
$\leq$ $\mathbb{E}\left[\left\|\hat{f}_{n-1}-\gamma_{n}\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)K_{L_{n}}^{T}(F(X_{n}),\cdot)-f^{*}\right\|_{K}^{2}\right]$
$\leq$ $\mathbb{E}\left[\left\|\hat{f}_{n-1}-f^{*}\right\|_{K}^{2}\right]-2\gamma_{n}\mathbb{E}\left[\left\langle\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)K_{L_{n}}^{T}(F(X_{n}),\cdot),\hat{f}_{n-1}-f^{*}\right\rangle_{K}\right]+\gamma_{n}^{2}M_{1}^{2}$
$\overset{\text{(i)}}{=}$ $\mathbb{E}\left[\left\|\hat{f}_{n-1}-f^{*}\right\|_{K}^{2}\right]-2\gamma_{n}\mathbb{E}\left[\left\langle\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)K_{L_{n}}^{T}(F(X_{n}),\cdot),\hat{f}_{n-1}-f_{L_{n}}\right\rangle_{K}\right]+\gamma_{n}^{2}M_{1}^{2},$
where (i) follows from the orthogonality between $K_{L_{n}}^{T}(F(X_{n}),\cdot)\in\mathcal{H}{L_{n}}$ and $f_{L_{n}}-f^{*}\in\mathcal{H}_{L_{n}}^{\perp}$. Next, we consider the second term in the final expression of (A.35)
$\mathbb{E}\left[\left\langle\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)K_{L_{n}}^{T}(F(X_{n}),\cdot),\hat{f}_{n-1}-f_{L_{n}}\right\rangle_{K}\right]$ (A.36)
$\overset{\text{(i)}}{=}$ $\mathbb{E}\left[\left\langle\mathbb{E}\left[\partial_{u}\ell\left(\hat{f}_{n-1}\circ F(X_{n}),Y_{n}\right)K_{L_{n}}^{T}(F(X_{n}),\cdot)\big|\ \mathcal{D}_{n-1}\right],\hat{f}_{n-1}-f_{L_{n}}\right\rangle_{K}\right]$
$=$ $\mathbb{E}\left[\left\langle\nabla\mathcal{E}(\hat{f}_{n-1})\big|_{\mathcal{H}_{L_{n}}},\hat{f}_{n-1}-f_{L_{n}}\right\rangle_{K}\right]$
$\overset{\text{(ii)}}{\geq}$ $\mathbb{E}\left[\mathcal{E}(\hat{f}_{n-1})-\mathcal{E}(f_{L_{n}})\right].$
In (i), we define $\mathcal{D}_{n-1}$ as the $\sigma$-field generated by the observations
$\mathcal{D}_{n-1}=\sigma\left((X_{1},Y_{1}),\dots,(X_{n-1},Y_{n-1})\right).$
In (ii), we use the convexity of $\mathcal{E}(f)$ on the set $\mathcal{W}\cap\mathcal{H}_{L_{n}}$, as established in Lemma A.5.
Substituting (A.36) into (A.35) yields
$\mathbb{E}\left[\left\|\hat{f}_{n}-f^{*}\right\|_{K}^{2}\right]\leq\mathbb{E}\left[\left\|\hat{f}_{n-1}-f^{*}\right\|_{K}^{2}\right]-2\gamma_{n}\mathbb{E}\left[\mathcal{E}(\hat{f}_{n-1})-\mathcal{E}(f_{L_{n}})\right]+\gamma_{n}^{2}M_{1}^{2}$
$\Rightarrow$ $2\gamma_{n}\mathbb{E}\left[\mathcal{E}(\hat{f}_{n-1})-\mathcal{E}(f_{L_{n}})\right]\leq\mathbb{E}\left[\left\|\hat{f}_{n-1}-f^{*}\right\|_{K}^{2}\right]-\mathbb{E}\left[\left\|\hat{f}_{n}-f^{*}\right\|_{K}^{2}\right]+\gamma_{n}^{2}M_{1}^{2}$
$\Rightarrow$ $\mathbb{E}\left[\mathcal{E}(\hat{f}_{n-1})-\mathcal{E}(f_{L_{n}})\right]\leq\frac{1}{2\gamma_{n}}\left(\mathbb{E}\left[\left\|\hat{f}_{n-1}-f^{*}\right\|_{K}^{2}\right]-\mathbb{E}\left[\left\|\hat{f}_{n}-f^{*}\right\|_{K}^{2}\right]\right)+\frac{\gamma_{n}}{2}M_{1}^{2}.$
Summing the above inequality from $(1-\alpha)n+1$ to $n$, we obtain
$\sum_{k=(1-\alpha)n+1}^{n}\mathbb{E}\left[\mathcal{E}(\hat{f}_{k-1})-\mathcal{E}(f_{L_{k}})\right]$
$\leq$ $\sum_{k=(1-\alpha)n+1}^{n}\frac{1}{2\gamma_{k}}\left(\mathbb{E}\left[\left\|\hat{f}_{k-1}-f^{*}\right\|_{K}^{2}\right]-\mathbb{E}\left[\left\|\hat{f}_{k}-f^{*}\right\|_{K}^{2}\right]\right)+\sum_{k=(1-\alpha)n+1}^{n}\frac{\gamma_{k}}{2}M_{1}^{2}$
$\leq$ $\frac{1}{2\gamma_{(1-\alpha)n}}\mathbb{E}\left[\left\|\hat{f}_{(1-\alpha)n}-f^{*}\right\|_{K}^{2}\right]$
$+\sum_{k=(1-\alpha)n}^{n-1}\mathbb{E}\left[\left\|\hat{f}_{k}-f^{*}\right\|_{K}^{2}\right]\left(\frac{1}{2\gamma_{k+1}}-\frac{1}{2\gamma_{k}}\right)+\sum_{k=(1-\alpha)n+1}^{n}\frac{\gamma_{k}}{2}M_{1}^{2}$
$\overset{\text{(i)}}{\leq}$ $\left[\frac{\widetilde{C}}{2\gamma_{0}}+\frac{2r\widetilde{C}}{\gamma_{0}}+\frac{\gamma_{0}}{2}M_{1}^{2}(2r+1)\right]\log(n+1)n^{\frac{1}{2r+1}},$
Here, we obtain (i) by applying the estimate from Lemma A.15. By Jensen’s inequality for the convex function $\mathcal{E}(f)$ on $\mathcal{W}$, we have
$\mathbb{E}\left[\mathcal{E}\left(\bar{f}_{\alpha n}\right)-\frac{1}{\alpha n}\sum_{k=(1-\alpha)n+1}^{n}\mathcal{E}\left(f_{L_{k}}\right)\right]\leq\frac{1}{\alpha n}\sum_{k=(1-\alpha)n+1}^{n}\mathbb{E}\left[\mathcal{E}(\hat{f}_{k-1})-\mathcal{E}(f_{L_{k}})\right]$ (A.37)
$\leq$ $\frac{1}{\alpha}\left[\frac{\widetilde{C}}{2\gamma_{0}}+\frac{2r\widetilde{C}}{\gamma_{0}}+\frac{\gamma_{0}}{2}M_{1}^{2}(2r+1)\right]\log(n+1)n^{-\frac{2r}{2r+1}}.$
Then we consider to bound the term
$\frac{1}{\alpha n}\sum_{k=(1-\alpha)n+1}^{n}\left[\mathcal{E}\left(f_{L_{k}}\right)-\mathcal{E}(f^{*})\right]\overset{\text{(i)}}{\leq}\frac{1}{\alpha n}\sum_{k=(1-\alpha)n+1}^{n}\frac{L}{2}\left\|f_{L_{k}}\circ F-f^{*}\circ F\right\|_{\rho_{X}}^{2}$ (A.38)
$\overset{\text{(ii)}}{\leq}$ $\frac{1}{\alpha n}\frac{L}{2}B_{\rho}\Omega_{d-1}A_{1}^{2r}\|g^{*}\|_{\omega}^{2}\sum_{k=(1-\alpha)n+1}^{n}(k+1)^{-\frac{2r}{2r+1}}$
$\leq$ $\frac{1}{\alpha n}\frac{L}{2}B_{\rho}\Omega_{d-1}A_{1}^{2r}\|g^{*}\|_{\omega}^{2}\int_{x=(1-\alpha)n}^{n}x^{-\frac{2r}{2r+1}}dx$
$\leq$ $\frac{(2r+1)LB_{\rho}\Omega_{d-1}A_{1}^{2r}\|g^{*}\|_{\omega}^{2}}{2\alpha}n^{-\frac{2r}{2r+1}}.$
In (i), we apply Lemma A.3, and in (ii), we apply Lemma A.12. Finally, we complete the proof by combining (A.37) and (A.38).
$\mathbb{E}\left[\mathcal{E}\left(\bar{f}_{\alpha n}\right)-\mathcal{E}\left(f^{*}\right)\right]$
$\leq$ $\mathbb{E}\left[\mathcal{E}\left(\bar{f}_{\alpha n}\right)-\frac{1}{\alpha n}\sum_{k=(1-\alpha)n+1}^{n}\mathcal{E}\left(f_{L_{k}}\right)\right]+\frac{1}{\alpha n}\sum_{k=(1-\alpha)n+1}^{n}\left[\mathcal{E}\left(f_{L_{k}}\right)-\mathcal{E}(f^{*})\right]$
$\leq$ $\frac{1}{\alpha}\left[\frac{\widetilde{C}}{2\gamma_{0}}+\frac{2r\widetilde{C}}{\gamma_{0}}+\frac{\gamma_{0}}{2}M_{1}^{2}(2r+1)+\frac{(2r+1)LB_{\rho}\Omega_{d-1}A_{1}^{2r}\|g^{*}\|_{\omega}^{2}}{2\log(2)}\right]\log(n+1)n^{-\frac{2r}{2r+1}}.$
A.3.2 Convergence Analysis of the Last Iteration
In this section, we use the results from Subsection A.2 and subsubsection A.3.1 to analyze the convergence of $\hat{f}_{n}$. First, we choose $0\leq m\leq i\leq n$, so that $\hat{f}_{i},\hat{f}_{m}\in\mathcal{H}_{L_{i}}\cap\mathcal{W}$, and we have
$\mathbb{E}\left[\left\|\hat{f}_{i+1}-\hat{f}_{m}\right\|_{K}^{2}\right]$
$=$ $\mathbb{E}\left[\left\|P_{\mathcal{W}}\left(\hat{f}_{i}-\gamma_{i+1}\partial_{u}\ell\left(\hat{f}_{i}\circ F(X_{i+1}),Y_{i+1}\right)K_{L_{i+1}}^{T}\left(F(X_{i+1}),\cdot\right)\right)-\hat{f}_{m}\right\|_{K}^{2}\right]$
$\leq$ $\mathbb{E}\left[\left\|\hat{f}_{i}-\gamma_{i+1}\partial_{u}\ell\left(\hat{f}_{i}\circ F(X_{i+1}),Y_{i+1}\right)K_{L_{i+1}}^{T}\left(F(X_{i+1}),\cdot\right)-\hat{f}_{m}\right\|_{K}^{2}\right]$
$\leq$ $\mathbb{E}\left[\left\|\hat{f}_{i}-\hat{f}_{m}\right\|_{K}^{2}\right]-2\gamma_{i+1}\mathbb{E}\left[\left\langle\partial_{u}\ell\left(\hat{f}_{i}\circ F(X_{i+1}),Y_{i+1}\right)K_{L_{i+1}}^{T}\left(F(X_{i+1}),\cdot\right),\hat{f}_{i}-\hat{f}_{m}\right\rangle_{K}\right]+\gamma_{i+1}^{2}M_{1}^{2}.$
Since $\mathcal{E}(f)$ is convex on $\mathcal{W}$, we have
$\mathbb{E}\left[\mathcal{E}\left(\hat{f}_{i}\right)-\mathcal{E}\left(\hat{f}_{m}\right)\right]$ (A.39)
$\leq$ $\mathbb{E}\left[\left\langle\partial_{u}\ell\left(\hat{f}_{i}\circ F(X_{i+1}),Y_{i+1}\right)K_{L_{i+1}}^{T}\left(F(X_{i+1}),\cdot\right),\hat{f}_{i}-\hat{f}_{m}\right\rangle_{K}\right]$
$\leq$ $\frac{1}{2\gamma_{i+1}}\left(\mathbb{E}\left[\left\|\hat{f}_{i}-\hat{f}_{m}\right\|_{K}^{2}\right]-\mathbb{E}\left[\left\|\hat{f}_{i+1}-\hat{f}_{m}\right\|_{K}^{2}\right]\right)+\frac{\gamma_{i+1}}{2}M_{1}^{2},$
We sum both sides of (A.39) from i=nki=n-k to $n$, where $k$ is an integer such that $1\leq k\leq\frac{n}{2}$, and set m=nkm=n-k
$\sum_{i=n-k}^{n}\mathbb{E}\left[\mathcal{E}\left(\hat{f}_{i}\right)-\mathcal{E}\left(\hat{f}_{n-k}\right)\right]$ (A.40)
$\leq$ $\sum_{i=n-k}^{n}\frac{1}{2\gamma_{i+1}}\left(\mathbb{E}\left[\left\|\hat{f}_{i}-\hat{f}_{n-k}\right\|_{K}^{2}\right]-\mathbb{E}\left[\left\|\hat{f}_{i+1}-\hat{f}_{n-k}\right\|_{K}^{2}\right]\right)+\sum_{i=n-k}^{n}\frac{\gamma_{i+1}}{2}M_{1}^{2}$
$\leq$ $\sum_{i=n-k+1}^{n}\mathbb{E}\left[\left\|\hat{f}_{i}-\hat{f}_{n-k}\right\|_{K}^{2}\right]\left(\frac{1}{2\gamma_{i+1}}-\frac{1}{2\gamma_{i}}\right)+\sum_{i=n-k}^{n}\frac{\gamma_{i+1}}{2}M_{1}^{2}$
$\overset{\text{(i)}}{\leq}$ $\left[\frac{8\widetilde{C}}{\gamma_{0}}+\gamma_{0}M_{1}^{2}\right](k+1)(n+1)^{-\frac{2r}{2r+1}}\log(n+2),$
where (i) is due to Lemma A.17.
Let $S_{k}=\frac{1}{k+1}\sum_{i=n-k}^{n}\mathbb{E}\left[\mathcal{E}\left(\hat{f}_{i}\right)\right]$ denote the average expected population risk over the last $k+1$ iterations. Then, by applying (A.40), we obtain
$\mathbb{E}\left[\mathcal{E}\left(\hat{f}_{n-k}\right)\right]\leq-S_{k}+\left[\frac{8\widetilde{C}}{\gamma_{0}}+\gamma_{0}M_{1}^{2}\right](n+1)^{-\frac{2r}{2r+1}}\log(n+2).$ (A.41)
Combining the definition of $S_{k}$ with (A.41) yields
$kS_{k-1}$ $=(k+1)S_{k}-\mathbb{E}\left[\mathcal{E}\left(\hat{f}_{n-k}\right)\right]=kS_{k}+\left(S_{k}-\mathbb{E}\left[\mathcal{E}\left(\hat{f}_{n-k}\right)\right]\right)$ (A.42)
$\leq kS_{k}+\left[\frac{8\widetilde{C}}{\gamma_{0}}+\gamma_{0}M_{1}^{2}\right](n+1)^{-\frac{2r}{2r+1}}\log(n+2)$
$\Rightarrow\quad S_{k-1}$ $\leq S_{k}+\frac{1}{k}\left[\frac{8\widetilde{C}}{\gamma_{0}}+\gamma_{0}M_{1}^{2}\right](n+1)^{-\frac{2r}{2r+1}}\log(n+2).$
Applying (A.42) recursively for $k=0$ to $\frac{n}{2}$, we obtain
$\mathbb{E}\left[\mathcal{E}\left(\hat{f}_{n}\right)\right]$ $=S_{0}\leq S_{\frac{n}{2}}+\left[\frac{8\widetilde{C}}{\gamma_{0}}+\gamma_{0}M_{1}^{2}\right](n+1)^{-\frac{2r}{2r+1}}\log(n+2)\sum_{k=1}^{\frac{n}{2}}\frac{1}{k}$
$\leq S_{\frac{n}{2}}+\left[\frac{8\widetilde{C}}{\gamma_{0}}+\gamma_{0}M_{1}^{2}\right](n+1)^{-\frac{2r}{2r+1}}\log(n+2)\left(1+\log\left(\frac{n}{2}\right)\right)$
$\leq S_{\frac{n}{2}}+2\left[\frac{8\widetilde{C}}{\gamma_{0}}+\gamma_{0}M_{1}^{2}\right](n+1)^{-\frac{2r}{2r+1}}\left(\log(n+2)\right)^{2}.$
Based on the estimates of inequalities (A.37) and (A.38) in the convergence analysis of $\alpha$-suffix averaging, we obtain
$S_{\frac{n}{2}}-\mathcal{E}(f^{*})$
$\leq$ $2\left[\frac{\widetilde{C}}{2\gamma_{0}}+\frac{2r\widetilde{C}}{\gamma_{0}}+\frac{\gamma_{0}}{2}M_{1}^{2}(2r+1)+\frac{(2r+1)LB_{\rho}\Omega_{d-1}A_{1}^{2r}\|g^{*}\|_{\omega}^{2}}{2\log(2)}\right]\log(n+1)n^{-\frac{2r}{2r+1}}.$
Combining the two estimates above, we obtain the error bound for the last iteration stated in the Theorem 1,
$\mathbb{E}\left[\mathcal{E}\left(\hat{f}_{n}\right)-\mathcal{E}(f^{*})\right]$
$\leq$ $2\left[\frac{8\widetilde{C}}{\gamma_{0}}+\gamma_{0}M_{1}^{2}\right](n+1)^{-\frac{2r}{2r+1}}\left(\log(n+2)\right)^{2}$
$+2\left[\frac{\widetilde{C}}{2\gamma_{0}}+\frac{2r\widetilde{C}}{\gamma_{0}}+\frac{\gamma_{0}}{2}M_{1}^{2}(2r+1)+\frac{(2r+1)LB_{\rho}\Omega_{d-1}A_{1}^{2r}\|g^{*}\|_{\omega}^{2}}{2\log(2)}\right]\log(n+1)n^{-\frac{2r}{2r+1}}.$
A.3.3 Technical Results
Lemma A.15.
Assuming that the assumptions and conclusions of Theorem 2 hold, then we have
$\frac{1}{2\gamma_{(1-\alpha)n}}\mathbb{E}\left[\left\|\hat{f}_{(1-\alpha)n}-f^{*}\right\|_{K}^{2}\right]$
$+\sum_{k=(1-\alpha)n}^{n-1}\mathbb{E}\left[\left\|\hat{f}_{k}-f^{*}\right\|_{K}^{2}\right]\left(\frac{1}{2\gamma_{k+1}}-\frac{1}{2\gamma_{k}}\right)+\sum_{k=(1-\alpha)n+1}^{n}\frac{\gamma_{k}}{2}M_{1}^{2}$
$\leq$ $\left[\frac{\widetilde{C}}{2\gamma_{0}}+\frac{2r\widetilde{C}}{\gamma_{0}}+\frac{\gamma_{0}}{2}M_{1}^{2}(2r+1)\right]\log(n+1)n^{\frac{1}{2r+1}}.$
Proof.
We now present the proof directly
$\frac{1}{2\gamma_{(1-\alpha)n}}\mathbb{E}\left[\left\|\hat{f}_{(1-\alpha)n}-f^{*}\right\|_{K}^{2}\right]$
$+\sum_{k=(1-\alpha)n}^{n-1}\mathbb{E}\left[\left\|\hat{f}_{k}-f^{*}\right\|_{K}^{2}\right]\left(\frac{1}{2\gamma_{k+1}}-\frac{1}{2\gamma_{k}}\right)+\sum_{k=(1-\alpha)n+1}^{n}\frac{\gamma_{k}}{2}M_{1}^{2}$
$\leq$ $\frac{\left((1-\alpha)n\right)^{\frac{2r}{2r+1}}}{2\gamma_{0}\log\left((1-\alpha)n+1\right)}\widetilde{C}\left(\log\left((1-\alpha)n+1\right)\right)^{2}\left((1-\alpha)n+1\right)^{-\frac{2r-1}{2r+1}}$
$+\frac{\widetilde{C}}{2\gamma_{0}}\sum_{k=(1-\alpha)n}^{n-1}\left((\log(k+1)\right)^{2}(k+1)^{-\frac{2r-1}{2r+1}}\left(\frac{(k+1)^{\frac{2r}{2r+1}}}{\log(k+2)}-\frac{k^{\frac{2r}{2r+1}}}{\log(k+1)}\right)$
$+\frac{\gamma_{0}}{2}M_{1}^{2}\sum_{k=(1-\alpha)n+1}^{n}k^{-\frac{2r}{2r+1}}\log(k+1)$
$\leq$ $\frac{\widetilde{C}}{2\gamma_{0}}\log\left((1-\alpha)n+1\right)\left((1-\alpha)n\right)^{\frac{1}{2r+1}}$
$+\frac{\widetilde{C}}{2\gamma_{0}}\sum_{k=(1-\alpha)n}^{n-1}\left((\log(k+1)\right)^{2}(k+1)^{-\frac{2r-1}{2r+1}}\left(\frac{(k+1)^{\frac{2r}{2r+1}}}{\log(k+1)}-\frac{k^{\frac{2r}{2r+1}}}{\log(k+1)}\right)$
$+\frac{\gamma_{0}}{2}M_{1}^{2}\log(n+1)\sum_{k=(1-\alpha)n+1}^{n}k^{-\frac{2r}{2r+1}}$
$\leq$ $\frac{\widetilde{C}}{2\gamma_{0}}\log\left(n+1\right)n^{\frac{1}{2r+1}}+\frac{\widetilde{C}}{2\gamma_{0}}\log(n+1)\sum_{k=(1-\alpha)n}^{n-1}(k+1)^{-\frac{2r-1}{2r+1}}\left((k+1)^{\frac{2r}{2r+1}}-k^{\frac{2r}{2r+1}}\right)$
$+\frac{\gamma_{0}}{2}M_{1}^{2}\log(n+1)\int_{x=(1-\alpha)n}^{n}x^{-\frac{2r}{2r+1}}dx$
$\overset{\text{(i)}}{\leq}$ $\frac{\widetilde{C}}{2\gamma_{0}}\log\left(n+1\right)n^{\frac{1}{2r+1}}+\frac{\widetilde{C}}{2\gamma_{0}}\log(n+1)\sum_{k=(1-\alpha)n}^{n-1}(k+1)^{-\frac{2r-1}{2r+1}}\left(\frac{2r}{2r+1}k^{-\frac{1}{2r+1}}\right)$
$+\frac{\gamma_{0}}{2}M_{1}^{2}(2r+1)\log(n+1)n^{\frac{1}{2r+1}}$
$\overset{\text{(ii)}}{\leq}$ $\frac{\widetilde{C}}{2\gamma_{0}}\log\left(n+1\right)n^{\frac{1}{2r+1}}+\frac{2\widetilde{C}}{2\gamma_{0}}\log(n+1)\frac{2r}{2r+1}\sum_{k=(1-\alpha)n}^{n-1}(k+1)^{-\frac{2r}{2r+1}}$
$+\frac{\gamma_{0}}{2}M_{1}^{2}(2r+1)\log(n+1)n^{\frac{1}{2r+1}}$
$\leq$ $\frac{\widetilde{C}}{2\gamma_{0}}\log\left(n+1\right)n^{\frac{1}{2r+1}}+\frac{\widetilde{C}}{\gamma_{0}}(2r)\log(n+1)n^{\frac{1}{2r+1}}+\frac{\gamma_{0}}{2}M_{1}^{2}(2r+1)\log(n+1)n^{\frac{1}{2r+1}}$
$\leq$ $\left[\frac{\widetilde{C}}{2\gamma_{0}}+\frac{2r\widetilde{C}}{\gamma_{0}}+\frac{\gamma_{0}}{2}M_{1}^{2}(2r+1)\right]\log(n+1)n^{\frac{1}{2r+1}},$
where (i) follows from Lagrange’s mean value theorem. In (ii), we use the inequality $(k+1)^{\frac{1}{2r+1}}/k^{\frac{1}{2r+1}}\leq 2$. This completes the proof. ∎
Lemma A.16.
Assuming the conditions of Theorem 2 hold, then for $\frac{n}{2}\leq n-k\leq i\leq n$, we have
$\mathbb{E}\left[\left\|\hat{f}_{i}-\hat{f}_{n-k}\right\|_{K}^{2}\right]\leq 8\widetilde{C}\left(\log(i+1)\right)^{2}\left(n+1\right)^{-\frac{2r-1}{2r+1}}.$
Proof.
We complete the proof directly through the following derivation
$\mathbb{E}\left[\left\|\hat{f}_{i}-\hat{f}_{n-k}\right\|_{K}^{2}\right]$ $\leq 2\mathbb{E}\left[\left\|\hat{f}_{i}-f^{*}\right\|_{K}^{2}\right]+2\mathbb{E}\left[\left\|\hat{f}_{n-k}-f^{*}\right\|_{K}^{2}\right]$
$\leq 2\widetilde{C}\left(\log(i+1)\right)^{2}\left(i+1\right)^{-\frac{2r-1}{2r+1}}+2\widetilde{C}\left(\log(n-k+1)\right)^{2}\left(n-k+1\right)^{-\frac{2r-1}{2r+1}}$
$\leq 2\widetilde{C}\left(\log(i+1)\right)^{2}\left[\left(i+1\right)^{-\frac{2r-1}{2r+1}}+\left(n-k+1\right)^{-\frac{2r-1}{2r+1}}\right]$
$\leq 4\widetilde{C}\left(\log(i+1)\right)^{2}\left(\frac{n+1}{2}\right)^{-\frac{2r-1}{2r+1}}\leq 8\widetilde{C}\left(\log(i+1)\right)^{2}\left(n+1\right)^{-\frac{2r-1}{2r+1}}.$
Lemma A.17.
Assuming the conditions of Theorem 2 hold, and noting that the first term in the following inequality is defined in (A.40), we obtain
$\sum_{i=n-k+1}^{n}\mathbb{E}\left[\left\|\hat{f}_{i}-\hat{f}_{n-k}\right\|_{K}^{2}\right]\left(\frac{1}{2\gamma_{i+1}}-\frac{1}{2\gamma_{i}}\right)+\sum_{i=n-k}^{n}\frac{\gamma_{i+1}}{2}M_{1}^{2}$
$\leq$ $\left[\frac{8\widetilde{C}}{\gamma_{0}}+\gamma_{0}M_{1}^{2}\right](k+1)(n+1)^{-\frac{2r}{2r+1}}\log(n+2).$
Proof.
This proof is similar to that of Lemma A.15. We present the proof directly
$\sum_{i=n-k+1}^{n}\mathbb{E}\left[\left\|\hat{f}_{i}-\hat{f}_{n-k}\right\|_{K}^{2}\right]\left(\frac{1}{2\gamma_{i+1}}-\frac{1}{2\gamma_{i}}\right)+\sum_{i=n-k}^{n}\frac{\gamma_{i+1}}{2}M_{1}^{2}$
$\overset{\text{(i)}}{\leq}$ $\frac{8\widetilde{C}}{2\gamma_{0}}(n+1)^{-\frac{2r-1}{2r+1}}\sum_{i=n-k+1}^{n}\left(\log(i+1)\right)^{2}\left(\frac{(i+1)^{\frac{2r}{2r+1}}}{\log(i+2)}-\frac{i^{\frac{2r}{2r+1}}}{\log(i+1)}\right)$
$+\frac{\gamma_{0}M_{1}^{2}}{2}\sum_{i=n-k}^{n}(i+1)^{-\frac{2r}{2r+1}}\log(i+2)$
$\overset{\text{(ii)}}{\leq}$ $\frac{8\widetilde{C}}{2\gamma_{0}}(n+1)^{-\frac{2r-1}{2r+1}}\log(n+1)\sum_{i=n-k+1}^{n}i^{-\frac{1}{2r+1}}+\frac{\gamma_{0}M_{1}^{2}}{2}\sum_{i=n-k}^{n}(i+1)^{-\frac{2r}{2r+1}}\log(i+2)$
$\overset{\text{(iii)}}{\leq}$ $\frac{8\widetilde{C}}{2\gamma_{0}}k(n+1)^{-\frac{2r-1}{2r+1}}\log(n+1)\left(\frac{n+1}{2}\right)^{-\frac{1}{2r+1}}+\frac{\gamma_{0}M_{1}^{2}}{2}(k+1)\left(\frac{n+1}{2}\right)^{-\frac{2r}{2r+1}}\log(n+2)$
$\leq$ $\left[\frac{8\widetilde{C}}{\gamma_{0}}+\gamma_{0}M_{1}^{2}\right](k+1)(n+1)^{-\frac{2r}{2r+1}}\log(n+2),$
where (i) is due to the inequality in Lemma A.16:
$\mathbb{E}\left[\left\|\hat{f}_{i}-\hat{f}_{n-k}\right\|_{K}^{2}\right]\leq 8\widetilde{C}\left(\log(i+1)\right)^{2}\left(n+1\right)^{-\frac{2r-1}{2r+1}}.$
In (ii), we apply Lagrange’s mean value theorem and use the inequality $\frac{1}{\log(i+2)}\leq\frac{1}{\log(i+1)}$. In (iii), we use the condition $\frac{n}{2}\leq n-k\leq n$. This completes the proof. ∎
A.4 Proof of Proposition 1
In this section, we prove Proposition 1. By Euler’s inequality (Section 7.12-3 in [15]), we have for any $f\in\mathcal{W}$ that
$\left\langle\nabla\mathcal{E}(f^{*})\big|_{\mathcal{H}_{K}},f-f^{*}\right\rangle_{K}\geq 0.$
Combining this with the identity in Lemma A.5, we obtain
$\mathcal{E}(f)-\mathcal{E}(f^{*})\geq\frac{\mu}{2}\|f\circ F-f^{*}\circ F\|_{\rho_{X}}^{2}.$
Finally, we complete the proof by applying the following inequalities
$\mathbb{E}\left[\left\|\hat{f}_{n}\circ F-f^{*}\circ F\right\|_{\rho_{X}}^{2}\right]$ $\leq\frac{2}{\mu}\mathbb{E}\left[\mathcal{E}\left(\hat{f}_{n}\right)-\mathcal{E}\left(f^{*}\right)\right]\leq\mathcal{O}\left(n^{-\frac{2r}{2r+1}}\left(\log(n+1)\right)^{2}\right)$
$\mathbb{E}\left[\left\|\bar{f}_{\alpha n}\circ F-f^{*}\circ F\right\|_{\rho_{X}}^{2}\right]$ $\leq\frac{2}{\mu}\mathbb{E}\left[\mathcal{E}\left(\bar{f}_{\alpha n}\right)-\mathcal{E}\left(f^{*}\right)\right]\leq\mathcal{O}\left(n^{-\frac{2r}{2r+1}}\log(n+1)\right).$
A.5 Proof of Lemma 1
In this section, we provide the proof of Lemma 1. We consider the following Sobolev ellipsoid characterized by parameters $s>\frac{1}{2}$ and $r\geq\frac{1}{2}$, with $l:=\lim_{k\to\infty}a_{k}\cdot\left(\dim\Pi_{k}^{d}\right)^{2s}\in(0,\infty)$,
$\mathcal{S}(4sr,Q)=\left\{\sum_{k=0}^{\infty}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}f_{k,j}Y_{k,j}\,\big|\sum_{k=0}^{\infty}\sum_{j=1}^{\dim\mathcal{H}_{k}^{d}}\frac{f_{k,j}^{2}}{a_{k}^{2r}}\leq Q^{2}\right\}.$
[…]
$A_{2}^{r}d^{-2sr}\frac{1}{j^{2sr}}\leq\lambda_{j}\leq A_{1}^{r}\frac{1}{j^{2sr}}\quad\forall j\in{\mathbb{N}}.$
Using the rearranged orthonormal eigensystem ${(\lambda_{j},\phi_{j})}_{j\geq 1}$, the Sobolev ellipsoid $\mathcal{S}(4sr,Q)$ can be rewritten as
$\mathcal{S}(4sr,Q)=\left\{\sum_{j=1}^{\infty}f_{j}\phi_{j}\,\Bigg|\sum_{j=1}^{\infty}\frac{f_{j}^{2}}{\lambda_{j}^{2}}\leq Q^{2}\right\}.$
Analogous to the proof of Example 5.12 in [58], we obtain the asymptotic bounds for the metric entropy of $\mathcal{S}(4sr,Q)$. Specifically, there exist constants $A_{3}\geq 1\geq A_{4}>0$ such that
$A_{4}\left(\frac{1}{\delta}\right)^{\frac{1}{2sr}}\leq\log N\left(\delta;\mathcal{S}(4sr,Q),\|\cdot\|_{\omega}\right)\leq A_{3}\left(\frac{1}{\delta}\right)^{\frac{1}{2sr}}\quad\text{for all small enough }\delta>0.$
Here we take an arbitrary estimator $G_{n}=D_{n}\circ E_{n}$, which is an $l_{n}$-sized estimator as described in the theorem with $l_{n}=o\left(n^{\tfrac{1}{2s(2r+1)}}\right)$. We next introduce the notion of an $\epsilon$-net with respect to the decoder $D_{n}$, which is used to characterize the collection of $l_{n}$-sized estimators $G(l_{n})$ can approximate the function class under an error tolerance $\epsilon$,
$\text{net}\left(\epsilon,l_{n},D_{n},\mathcal{S}(4sr,Q)\right)=\left\{f\in\mathcal{S}(4sr,Q)\,\bigg|\,\exists\,b_{n}\in\{0,1\}^{l_{n}},\text{such that }\|f-D_{n}(b_{n})\|_{\omega}\leq\epsilon\right\}.$
Furthermore, by the definition of $l_{n}$, there exists a sequence $m_{n}$ such that $l_{n}=o(m_{n})$ and $m_{n}=o\left(n^{\frac{1}{2s(2r+1)}}\right)$. Here, setting $\delta=m_{n}^{-2sr}$, the metric entropy satisfies
$\log_{2}N\left(m_{n}^{-2sr};\mathcal{S}(4sr,Q),\|\cdot\|_{\omega}\right)\geq A_{4}\log_{2}(e)m_{n}\geq A_{4}m_{n}.$
Since $l_{n}=o(m_{n})$, the set $D_{n}\left(\{0,1\}^{l_{n}}\right)$, which contains at most $2^{l_{n}}$ elements, cannot form an $m_{n}^{-2sr}$-cover of $\mathcal{S}(4sr,Q)$ for sufficiently large $n$, namely
[…]
Let us denote $\alpha_{n}=E_{n}\left(\{(X_{i},Y_{i})\}_{1\leq i\leq n}\right)\in\{0,1\}^{l_{n}}$, one has
[…]
$\geq$ $\sup_{f^{*}\in\mathcal{S}(4sr,Q)}\mathbb{E}\left[\|G_{n}\left(\{(X_{i},Y_{i})\}_{1\leq i\leq n}\right)-f^{*}\|_{\omega}^{2}\right]$
$=$ $\sup_{f^{*}\in\mathcal{S}(4sr,Q)}\mathbb{E}\left[\|D_{n}\left(\alpha_{n}\right)-f^{*}\|_{\omega}^{2}\right]$
[…]
Consequently, we obtain
[…]
Taking the limit as $n\to\infty$ on both sides yields the conclusion of Lemma 1.